### v2

# Measuring Compliance of Consent Revocation on the Web

Abstract
The GDPR requires websites to facilitate the right to revoke consent from Web users. While numerous studies measured compliance of consent with the various consent requirements, no prior work has studied consent revocation on the Web. Therefore, it remains unclear how difficult it is to revoke consent on the websites’ interfaces, nor whether revoked consent is properly stored and communicated behind the user interface.
Our work aims to fill this gap by measuring compliance of consent revocation on the Web on the top-200 websites. We found that 19.87% of websites make it difficult for users to revoke consent throughout different interfaces, 20.5% of websites require more effort than acceptance, and 2.48% do not provide consent revocation at all, thus violating legal requirements for valid consent. 57.5% websites do not delete the cookies after consent revocation enabling continuous illegal processing of users’ data.
Moreover, we analyzed 281 websites implementing the IAB Europe TCF, and found 22 websites that store a positive consent despite user’s revocation. Surprisingly, we found that on 101 websites, third parties that have received consent upon user’s acceptance, are not informed of user’s revocation, leading to the illegal processing of users’ data by such third parties. Our findings emphasise the need for improved legal compliance of consent revocation, and proper, consistent, and uniform implementation of revocation communication and data deletion practices.
1 Introduction
In the past few years, compliance with privacy and data protection laws on the Web has gained a lot of attention, both from the research community and regulators. The EU Data Protection framework, consisting of the General Data Protection Regulation (GDPR) [1] and ePrivacy Directive (ePD) [2] sets requirements for a legally-valid consent when tracking technologies are deployed on a website or app. Such consent is usually implemented via consent banners and must satisfy seven requirements for validity: it must be prior to any data collection, freely given, specific, informed, unambiguous, readable and accessible, and revocable [1, Art. 4(11), 7] [3]. While numerous studies have evaluated compliance of consent banners with the various consent requirements, such as the presence and visibility of the rejection button  [4, 5, 6], one aspect has so far been overlooked by the research community and the regulators: the requirement of revocable consent.
According to the GDPR, users have the right to revoke consent at any time [1, Art. 7(3), Recital 42]; thus, websites must facilitate the exercise of this right [1, Art. 12(2)] by providing an option to revoke consent and delete related data. Websites that do not allow consent revocation are deemed to be processing data illegally, and run the risk of being fined for not complying with the legal requirements [7, 8, 9, 10, 11]. However, it remains unclear whether websites provide users with compliant methods to revoke their consent, whether revoked consent is properly stored by the websites behind the consent interface, and whether it is communicated to and deleted by third-parties that collected users’ data. Our work aims to fill this gap by measuring the compliance of consent revocation on the Web. We address the following questions:
  • RQ1:
    Are revocation interfaces compliant with EU data protection laws?
  • RQ2:
    Do advertising and analytics (AA) cookies that require consent get deleted once consent is revoked?
  • RQ3:
    Is consent stored correctly in the browser, and is it consistent across browser storage and APIs implemented by Consent Management Platforms (CMPs)?
  • RQ4:
    Are all third parties that are initially informed of “acceptance” also notified when the consent is revoked?
To address these questions, we have set up a team of computer scientists and a legal expert to measure compliance of consent revocation on the Web, making the following contributions:
  1. 1.
    We provide the first in-depth legal analysis of the GDPR, ePD and other legal guidelines for consent revocation interface, storage and communication, and establish six operational legal requirements (§3).
  2. 2.
    We provide a method to analyse consent revocation interfaces (§4.1), and apply our analysis on the top 200 websites identifying multiple instances of non-compliance with the legal requirements (§5).
  3. 3.
    We study the effect of revoking advertising and analytics (AA) cookies, finding that on the majority of websites revocation does not decrease the number of AA cookies, thus violating the GDPR (§6).
  4. 4.
    We propose a methodology (§4.2) to evaluate revocation on websites that implement IAB Europe Transparency and Consent Framework (TCF) and the most popular CMP, OneTrust. We analyse 281 such websites and evaluate the storage of consent behind the consent banner interface (§7) and the communication of revocation (§8), thereby detecting multiple non-compliant practices.
  5. 5.
    Finally, we give recommendations to further improve consent revocation compliance (§9).
Table I presents the most prevalent (more than 10%) potential violations of the GDPR and ePD that we have detected while analysing both consent revocation interface, and consent stored and shared with third parties behind the user interface.
Violation Prevalence
Revocation via a different interface/medium (§5) 19.87% (32/161)
Two or more steps to revoke vs zero steps to accept (§5) 20.5% (33/161)
Data processing based on AA cookies is not stopped upon user’s consent revocation(§6) 57.5% (69/120)
Positive consent after revocation (TCF-based consent string) (§7) 16.17% (22/136)
Positive consent after revocation (OneTrust-specific consent string) (§7) 14.47% (22/152)
Third-parties informed of consent acceptance via HTTP requests, but not informed of consent revocation (§8) 74.2% (101/136)
TABLE I: Summary of most prominent potential violations of consent revocation implementation. In brackets we mark the number of websites with violations and the total number of websites analysed for each case.
2 Background and Related Works
In this section, we provide background on different techniques for consent management, and discuss prior works on compliance in the context of consent banners, opt-out and revocation.
Consent Interfaces and Compliance. Consent banners have become a common method for obtaining consent for online tracking. Recent works [3, 12, 13, 14, 15, 16] have examined consent banners following the implementation of the GDPR and the ePD while numerous others [17, 18, 19, 20, 21, 22, 23, 4, 6, 24, 25] explored the impact of privacy regulations on consent design, and how the websites and CMPs might violate these laws, e.g., by using dark patterns. Tools have been proposed to help users handle consent banners automatically [26, 27, 28, 29, 5, 30, 31, 32]. Several works [5, 14, 6, 33] focused on automated analysis of consent banners at scale.
IAB Europe TCF and Compliance. The current “de-facto” standard for the consent banners in the EU is the IAB Europe Transparency and Consent Framework (TCF) [34]. The current TCF v2.2 [35] defines the pre-defined purposes for processing personal data by Consent Management Platforms (CMPs). If also defines the format to store the user’s choice, called TCString, that includes: (a) list of enabled third-party vendors registered within the TCF; (b) list of enabled purposes among the pre-defined purposes [36]. The TCF standard does not specify which of the purposes require user consent. According to GDPR and ePD, only purposes that do not require user consent (and users’ explicit interaction with the CMP) may be enabled in TCString  by default. Matte et al. [37] analysed which purposes in TCF v1.1 and v2.0 require consent according to the GDPR and ePD; however, this analysis was not extended to TCF v2.2, which we perform in this paper.
Once a user has made their choice on the CMP interface, the CMP is required to store the corresponding TCString in the browser (however, the specific storage mechanism is not specified in TCF v2.2), or implement an API for third-parties to check the TCString. The TCString can be accessed via the __tcfapi function by third-parties in first-party context or via the JavaScript postMessage API to communicate with a special __tcfLocator iFrame. Following TCF v1.1 [38], TCString can also be shared in the outgoing HTTP requests via URL-based methods. However, TCF v2.2 does not specify any URL parameters to be used for this.
Prior works [4, 37, 39, 40] have focused on the analysis of IAB Europe TCF and its consent implementation. However, none of these works studied how CMPs manage consent revocation, which is one of the main objectives of our work.
Opt-out and Consent Revocation. While multiple studies detected and analysed the presence of a reject button on consent banner interfaces [4, 6, 5], none of the prior works evaluated the interfaces that allow revocation of a given consent within the EU data protection framework. Nevertheless, since the updates of the CPRA regulation in California, multiple studies have analysed the implementation of the right to opt out of selling or sharing user data within websites. Tran et al. [41] propose a methodology to measure compliance of opt-out links’ wordings automatically found on websites subject to CCPA and CPRA [42]. Liu et al. [43] measure the impact of opting out in the presence of advertisers and tracking mechanisms.
Aziz et al. [25] assessed the IAB CCPA Compliance Framework (analogous to IAB TCF in Europe) and detected that opt-out signals are not honored on websites. A similar latest work by Du et al. [44] detected violation regarding withdrawal interfaces in mobile applications.
Other studies evaluated user perceptions when interacting with opt-out mechanisms. As such, Habib et al. [45] evaluated the usability of data deletion and opt-out options related to email communication and targeted advertising. Habib et al.[46] evaluated how different instructions about revocation in the banner’s text impact users’ ability to find revocation options.
None of these works, however, studied how difficult it is for the user to revoke consent, and whether revocation interfaces comply with the EU laws. Additionally, no prior work has shed light on whether websites correctly store and manage revoked consent, and inform all the third-parties that were previously informed of consent acceptance.
3 Legal Compliance for Consent Revocation
EU Data Protection framework provides legal principles and requirements that can be applied to websites and specifically to consent revocation. The GDPR applies to the processing of personal data [47], and the ePrivacy Directive (ePD) [2] provides supplementary rules, particularly, for the use of tracking technologies [48]. Whenever cookies and other tracking technologies are stored and read from the user’s device, the ePD [2, Art. 5(3)] requires websites to request user’s consent when tracking is used for purposes that are not strictly necessary for the service provided, e.g., advertising [49, 50]. Some purposes are exempted of consent, e.g, functional purposes required for a website to operate [2, Recital 66]. The way to assess with certainty whether consent is required is to analyze the purpose of each tracker on a given website [50, 51].
The GDPR, as well as guidelines from the Data Protection Authorities (DPAs) and from the European Data Protection Board (EU advisory board, representing all EU DPAs) provide legal requirements for consent revocation. While guidelines are not legally-binding, they are part of the EU framework for data protection which we apply in this work to discern when revocation methods are compliant. A website can be held liable and fined if it fails to comply with the GDPR principles and requirements for valid consent, including revocation. Relevant GDPR principles are presented below and numbered with P, and legal requirements for consent revocation for websites are numbered with LR.
P1 (Fairness).
Websites must not process personal data in a unjustifiably detrimental, discriminatory, unexpected or misleading way [1, Art. 5(1)(a)][52, para 17].
P2 (Data protection by design).
Websites must implement technical and organisational measures, and safeguards efficiently to enable the exercise of the revocation right [53, para 68], [1, Art. 25(1)]; [53, para 70]. Revocation options should be provided in an objective and neutral way, avoiding any deceptive or manipulative language or design [54, 55, para 16]. If a website requires more effort to revoke than to give consent, it is deploying a dark pattern [55, para 30].
P3 (Accountability).
Websites must be able to demonstrate that revocation is effective, i.e., performed easily and effectively [1, Arts. 5(2), 24(1), Rec. 74].
4 Methodology
In this section, we describe how the data was collected and the analysis we performed to answer our research questions. We built a semi-automated crawler based on a Selenium-instrumented Chromium v122.0.6261.128. We used the crawler to collect data in April-June, 2024111The code and dataset is available as supplementary material, see 10. within the EU. Table II presents an overview of the three datasets we collected, which are further explained in this section.
Figure 1: Data collection pipeline: to address RQ1, we collect screenshots and label a website as shown in violet boxes; for RQ2 we collect cookies at each stage and also at Rejection stage in a similar way (not shown in this figure).
4.1 Datasets for RQ1 and RQ2
We have implemented the data collection pipeline shown in Figure 1 to capture the elements of the user interface that need to be interacted with in order to exercise consent revocation. At each stage of our analysis, we save background data to further analyse cookies, consent storage and communication. Figure 2 presents the details of the background data collection.
Domains Collected dataset RQs
200
(ranked 1-200)
DS1: 105 domains without CMPs RQ1, RQ2
DS2: 56 domains with CMPs RQ1, RQ2,
RQ3, RQ4
1000 random
(ranked 201-5k)
DS3: 225 domains with CMPs RQ3, RQ4
TABLE II: Datasets used for RQs.
4.1.1 Website selection
We used the top-200 domains from Tranco [64, 65] for analysing the interface of revocation options. 161 domains were reachable while 39 domains did not display a webpage because they were either CDNs, were blocked or returned failures. We detected the CMP by decoding the TCString returned by the __tcfapi following the methodology of Matte et al. [4] after accepting consent. We additionally detect the websites using OneTrust CMP by the presence of OptanonConsent cookie, specific for this CMP. Of the 161 domains, 56 used TCF and/or OneTrust CMP (labeled “with CMP”), which we could detect.
4.1.2 Analyzing interfaces for consent revocation (RQ1)
We collected the data in a semi-automated manner as shown in Figure 1: we save screenshots to record the user interfaces involved in reaching a revocation option. While recent works [66, 67, 68] provide automated tools to select initial consent options on websites (specifically, button and toggle HTML elements) based on user choices, our analysis shows that revoking a previously given consent may require navigating multiple links on the page to reset the options once the initial banner is closed. Moreover, it is not always clear where the consent modification options are located on the website, and whether such options are present. Therefore, revocation options are difficult to find and interact with automatically due to heterogeneity of revocation implementations, as we show in this paper. The data is collected in four stages:
(1) Initial landing: We visit each website afresh with storage cleared for all websites in the browser, and collect the screenshots showing the presence of a banner.
(2) After acceptance: We record whether a banner is present and if an option is provided to the user to accept consent, in which case we accept all cookies.
(3) After revocation: After accepting consent, we check if an icon for modifying consent is visible on the page. If not, we check if an option is available in the footer or elsewhere on the main page. Finally, we check if the option is available on a different page reachable from the main page. For each of the options, we record the number of steps needed for revocation. If no option is found, we record that revocation is not possible.
(4) After rejection: To compare the behavior of the website after initial rejection (opt-out) with that after the consent is explicitly revoked, we clean the browser and record the options to reject consent on the website. If a consent banner is present, we reject all cookies.
Amid all these steps, we also take screenshots, in an automated manner, at different stages by scrolling to the bottom of the pages or navigating to the settings in order to get the important information like keywords leading to revocation. Using these inputs and screenshots, we categorize the websites based on the revocation options, where such options are provided, and the number of steps required to revoke consent. We then map these categories to the legal requirements for GDPR and consent revocation (from §3) to identify and measure potential violations of the law.
4.1.3 Effect of revocation on AA cookies (RQ2)
In background, we collect cookies stored in the browser at each of the four stages. Since only some cookies require consent (see §3), we followed Bouhoula et al. [68] and consider that only cookies classified as Advertising or Analytics by CookieBlock [5] require user consent. If we detect AA cookies upon initial landing, after rejection or after revocation, we map it to legal requirements and identify a potential violation with the law. Similarly, we analyse whether the number of AA cookies increases after revocation w.r.t. other stages of our analysis.
4.2 Datasets for RQ3 and RQ4
In websites that allowed revocation, different methods were used to store the modified consent and communicate it over the network. To detect the presence of CMPs that implement the IAB TCF [34], and therefore use a standardised format for storing consent (Transparency and Consent String or TCString), prior works [4, 40] used the __tcfapi function provided by TCF. We, therefore, performed an automated preliminary study to test this detection method. We also identify the detection methods for the use of OneTrust, which is the most popular CMP on Tranco top-20k websites [69].
4.2.1 Preliminary study
We performed an automated crawl querying for __tcfapi, a function that must be provided to use TCF’s functionalities, on the top 200 websites in the EU region, according to Tranco list [65]. Out of 161 websites that were reachable, we found __tcfapi on only 32 (20%) of websites [^2] Similarly to prior work [4]. , we found that __tcfLocator is present on only 1 extra website where the __tcfapi is not present. We therefore do not test for this method in this work.. We therefore further explored other ways to detect CMPs by analysing the implementations of the most popular CMP [69], OneTrust. By examining websites with OneTrust and its documentation [70], we found that this CMP (1) stores user consent in a very specific format in a cookie named OptanonConsent, and (2) maintains a JavaScript variable, OneTrustActiveGroups (OTAG), which can be queried to get the current consent string. However, the format of the consent string accessible through these methods does not respect the official IAB Europe TCF TCString format.
4.2.2 Website selection
We use the methods developed in the preliminary study to detect websites that use TCF and OneTrust from top-200 websites, resulting in dataset DS2 that contained only 56 websites with CMPs. We randomly chose additional 1000 websites between the rank 200 and 5000, and detected 225 websites that implemented either TCF or OneTrust, resulting in a set of 281 websites (DS2 and DS3).
Figure 2: Data collection pipeline: to address RQ3 (validity and consistency of consent), the collected sets of consent strings are highlighted in orange, while for RQ4 (access and communication of consent to 3rd-parties), the sets of scripts parties involved in the communication of consent are highlighted in violet. The objects in green represent our methods for collecting the data, and the boxes in yellow are further described in Table III.
4.2.3 Data collection
Figure 2 shows the data collection pipeline for RQ3 and RQ4 that we describe next. The data is collected during all four stages of the experiment (highlighted as “Save background data” in Figure 1). We collect the consent strings (both TCString and consent variable specific to OneTrust) for further analysis of validity and consistency of consent storage (RQ3), and complement it with the information about scripts that access consent strings via APIs, scripts that initiate the requests and third-parties with whom consent is shared for the analysis of communication of consent (RQ4).
4.2.4 Collecting stored consent strings and third-parties that access them
The current TCF v2.2, does not define the specific location where the user consent should be stored. Therefore, we collect all cookies and localStorage objects’ values that match the TCString format. Additionally, we collect the OptanonConsent cookie that OneTrust uses to store a user’s consent. We also inject scripts in the webpage to query __tcfapi and OTAG to analyse consent strings that third-parties would receive if they query these APIs, and examine network logs to detect and store TCStrings shared in the requests and responses. All the collected consent strings are visualised in Figure 2 in orange color.
Third-parties can access consent strings via APIs or receive them via network requests. We capture (1) third-party scripts that fetch consent string via __tcfapi or OTAG; (2) HTTP requests that contain consent strings in TCString format (we do not capture OneTrust-specific strings in the traffic because they do not have a well defined format). We store all initiators and receivers of consent strings (see violet stores in Figure 2).
4.2.5 Detecting and extracting TCStrings in network logs
To detect the TCString in network logs, we look at different parts of the request/response pair. Previous works have extracted TCStrings only from specific URL parameters, such as gdpr_consent [4] following TCF v1.1. A follow-up work [40] checked for outgoing TCString in all URL parameters of HTTP requests and also in the HTTP cookie headers, since TCF v2.2 does not specify URL parameters to be used. When parsing URLs, we extract all query parameters, for example for a URL https://www.site.com/zzz.jpg?ISBN=XXX&UID=ABC123, we first extract the values XXX and ABC123 and then check whether they correspond to the TCString format, according to TCF specification [71]. The orange boxes in Figure 2 represent the functions we used to extract TCStrings from different parts of the network logs that are further explained in Table III.
Since TCStrings do not have any protection mechanisms, any third-party that receives a TCString, can potentially modify it and return a different TCString back to the browser. To detect inconsistencies in TCString between requests and responses, we analyse postData in POST requests, and the response data (HTML and JSON formats) from the servers. In each case, we record the request initiator to identify scripts that sent wrong consent strings. To analyse HTTP GET requests, we search for TCString in the URL parameters, while for POST methods, we extract the TCString from JSON object if the data is sent in this format. To analyse HTTP responses and detect when TCString is modified by the server, we analyse redirected URLs: if response type is JSON, we parse it; otherwise, we parse the returned HTML file to extract TCString from the script, image and iFrame elements. Our approach uncovers more instances of TCStrings as compared to previous works [4, 40], since we detect 4,445 TCStrings in URL parameters, and an additional 3,423 TCStrings in the postData, 52 TCStrings in JSON objects and 170 TCStrings in HTML files. We also record requests where the cookie containing TCStrings is sent to third-parties.
Function in Fig. 2 Extracting TCString
parseURL (URL) Extract all query parameters from URL and keep all parameter values that match a TCString format.
parseJSON (obj) Extract all {key:value} pairs present in obj and keep all values that match a TCString format.
parseHTML (file) Extract URLs for <img>, <iframe> and <script> in file and call parseURL to extract TCStrings.
TABLE III: Parsing the network logs for extracting TCStrings
4.2.6 Wrappers for __tcfapi and OTAG
To observe who is accessing the __tcfapi, we override the function to record all the scripts which accessed it. We also record what command was requested through the call, e.g., “ping”, “getTCData” or “getEventHandler”. A similar approach was adopted by Matte et al. [4], however, they observed calls made to the APIs in an older version of TCF and their analysis did not include OneTrust. We override getter and setter methods of the OTAG variable to record which domain is writing/updating it (mostly by a designated CMP script) and reading/accessing it.
4.2.7 Communicating consent modification
We check the network logs to investigate if third-parties that received the TCString after acceptance had also received the updated consent after revocation. We do this by observing requests to third-party URLs, and if the request contained the TCString either in the URL or in the POST data as described in the Figure 2.
4.2.8 Identifying responsible parties
We record the initiators of all the network requests, and use this information to identify which third-party script changed the consent. We match script’s domain of provenance with the CMP name decoded from the TCStrings. If the name and domain do not match, in case of inconsistencies, we manually check the domains for cases where the CMP names are either shortened or have words like ‘privacy center’ or ‘CDN’. If we do not identify any such indicator that the third-party could be the CMP, we classify the responsible party just as a third-party.
4.3 Limitations
4.3.1 Accuracy of AA cookie classification
Even though CookieBlock [5] has better accuracy than other tools for classifying cookie purposes, it might not be entirely accurate: a website may declare some of the cookies, classified as AA, as necessary in their cookie policies or banners. Such cookies, classified by CookieBlock, will be, nevertheless, labeled as AA and considered to require consent.
4.3.2 Not capturing network logs while searching for revocation option on secondary pages
When reaching the revocation settings takes multiple steps, for a few seconds, log-entries with the old consent (after acceptance) are also saved along with the network logs meant for “after revocation” stage. We remove these log-entries by searching for inconsistencies after the consent is revoked. However, the set of log-entries misclassified as “after revocation” are not checked for inconsistencies and we may have missed a few cases similar to the ones mentioned in Table VIII in the Appendix.
4.3.3 Unable to capture access to consent when the third-parties use event listeners
Callback functions returned as an event listener by __tcfapi does not have a standard format or name for us to track the access requests made to the API. Similarly, OneTrust suggests third-party ad-systems like Google Tag Manager to implement their own event listeners to be updated when the user changes the consent. Hence, we do not track API accesses when event listeners are used.
4.3.4 Determining description of purposes in OneTrust
We do not know how the purposes displayed by OneTrust or any CMP used by the website are mapped to the user’s actual consent. Since there is no particular format used to store the consent string, it can vary between different CMPs, and within websites using the same CMP, as well.
4.3.5 Delay in registering user’s consent
While collecting network logs to assess the communication of consent revocation to third parties, we observed (see Section .1) that some websites may take time to register that a user revoked their initial consent. To account for this, we exclude cases with delays in updating consent strings from our list of violations. However, we include cases where some third parties correctly include the updated consent string in their network requests, while others still send the old consent (positive consent string). Because the implementation for notifying third parties is server-side, this may result in some false positive violations on our end.
5 Revocation Interface and its Compliance
This section examines user interfaces for revoking consent to determine if they potentially violate EU legal requirements, addressing RQ1. We first categorize the $161161$ reachable websites out of top 200 (datasets DS1 and DS2 from Table II), based on their consent banners: 108 (67%) websites display a consent banner to the users on their first visit, 8 (5%) websites displayed a banner with no option, not allowing neither to accept nor reject consent, and 45 (28%) websites did not display any banner. [^3] This can be explained by the fact that these websites were hosted outside of the EU and did not yet implement compliance solutions with EU laws. We further manually classify 161 websites based on the interface provided to revoke consent: 120 (74.5%) out of 161 websites provide users option to revoke consent within the same interface, medium or website where the consent request took place, or navigates to related pages for revoking consent; 32 (19.8%) websites provided users options to withdraw consent via different interface, where the options to revoke consent are present outside the interface or medium from where the consent request took place, and 9 (5.6%) websites offer no revocation option.
Within the Same Interface Via Different Interface No Revocation
Banner #
Icon
Footer
Options
Banner
on Policy
Options
via Policy
Settings
or links
Contact/
Email
After
Login
No option
provided
Consent banner 108 9 63 6 21 8 0 1 0
No option banner 8 0 0 0 5 2 0 1 0
Without banner 45 0 4 5 7 17 3 0 4 [+5]
Total 161 9 67 11 33 27 3 2 9
TABLE IV: Prevalence and type of consent revocation options on 161 websites (rank 1-200) from DS1. We use different colors to represent the level of compliance of the detected practices. Green color represents zero or one steps, yellow represents two or more steps, and pink and red represent violations of varying severity.
Table IV summarizes the results for these three categories. To further evaluate legal compliance, for websites that provide revocation within the same interface, we count the number of steps required to revoke consent based on Figure 1. Figure 6 (in the Appendix) shows examples of these categories.
5.1 Results
5.1.1 Compliant revocation interface implementation (zero or one steps)
Out of the 161 websites, only 9 (5.6%) offer a persistent icon or button floating on the page, thus requiring zero steps to reach the revocation option. $6767$ (41.6%) websites offered a link option in the footer of the page, requiring one step to reach the revocation option. 11 (6.8%) websites showed a consent banner or icon when accessing the privacy policy page from the footer, therefore also requiring one step to revoke consent. All these implementations found on 87 (54%) websites (labeled as Icon, Footer Options and Banner on Policy in Table IV) comply with revocation GDPR principles and consent requirements (LR1-3) since they are presented within the same interface and require zero or one step to revoke consent. According to Habib et al. [46], users who face persistent icons are more likely to recognise a correct method to revoke consent with respect to users who saw a link in the website’s footer. Therefore, even though all such implementations are compliant, only 7.4% websites with an Icon provide a more usable revocation design.
5.1.2 Two or more steps to revoke vs zero steps to accept
33 (20.5%) websites out of 161 allow users to access the revocation option within the same interface, but with additional obstruction, requiring them 2 or more steps to revoke consent. Such websites (labeled with Options via Policy) hid the option to revoke consent inside the “Cookie Policy” or “Privacy Policy” page. According to the majority of EU regulators, this additional effort does not allow to exercise the right of revocation in a easy and effective way even if such option is located withing the interface. Arguably, the websites requiring additional effort neither comply with LR3 (Easy revocation through the same effort and number of steps) nor with the principles P1 (Fairness), since websites require unjustifiable and unexpected effort, nor P2 (Data Protection by Design), since the adopted measures are not efficient as to facilitate the revocation right but obstruct it instead, and P3 (Accountability), since such websites are not able to demonstrate compliance with the requirement LR3.
5.1.3 Revocation options Via Different Interface
Overall, 32 (19.87%) websites out of 161 offered the option to revoke consent via an interface that is substantially different from the interface for users to accept consent. Figure 7 in the Appendix shows examples of websites for each of the options given.
Twenty-five websites, including medium.com, discord.com and wikipedia.org, suggest revocation through browser settings by clearing cookies or by offering opt-out links for third-party tracking or advertisement domains (labeled as Settings or links). Three websites – github.io, archive.org and who.int – suggested users to contact or email them to revoke consent or delete the data (labeled Contact/email).
Settings or links as well as Contact/email revocation options infringe the legal requirement LR2 (Easy revocation through the same interface), since revoking consent is not made available through the same means or interface; and principles P1 and P2, since switching to these totally different interfaces requires unnecessary and disruptive effort that is obstructive and unexpected. This leaves users in an asymmetrical relationship between giving and revoking consent, which does not permit users to exercise their revocation right. Consequently, these websites cannot be accountable for demonstrating compliance with this right (P3).
Two websites’ (tumblr.com and baidu.com) privacy settings lead the user to the login page and instruct the user to revoke consent after logging-in, which may require account creation (labeled After login). While this option occurs through the same website, it directs the user to a different interface. As before, this option infringes the requirement LR2 since the option redirects users to an unrelated interface. It also infringes principles P1 and P2 as it forces users to take unexpected additional steps and effort – to subscribe, and thus to give more personal data – to revoke consent. This option impedes these websites to show accountable revocation (P3).
5.1.4 No revocation option provided
Nine (5.6%) websites out of 161 did not provide any means to revoke consent, though they also did not display any consent banner, indicating that these websites probably did not integrate online tracking. We further analysed these websites and found that 4 (2.48%) of them stored AA cookies making them non-complaint. These websites include ntp.org and un.org. All these 4 websites that use AA cookies infringe the legal requirement LR1 (Right to revoke consent) and the following principles: P1 since websites process data unjustifiably for AA purposes without a legal basis, or the knowledge or consent of users; P2 as it does not create measures to enable the exercise of the revocation right. Consequently, these websites cannot be accountable for demonstrating compliance with this right (P3).
Figure 3: Revocation methods on top 161 websites
5.2 Impact of CMPs on compliance
Figure 3 shows the prevalence of different revocation options on the websites with detected CMPs vis a vis websites where we did not detect any CMP. 43 (77%) out of 56 websites with detected CMPs provide a compliant implementation (zero or one step) to revoke consent, while only 32 (30.5%) websites out of 105 websites without detected CMP have compliant implementation, showing overall a higher compliance rate for websites with CMPs.
Regarding non-compliant implementation, 37% (39 out of 105) websites without CMP require two steps or more to revoke consent, while only 10.7% (6 out of 56) websites with CMP needed 2 or more steps for revocation. None of the websites with CMP denied a revocation option to users, though 8.5% (9 out of 105) websites without CMP offered no revocation option. In conclusion, websites with detected CMPs tend to be more compliant with revocation requirements.
6 Effect of Revocation on AA Cookies
(a) Comparison of AA cookies upon
initial landing vs after revocation
(b) Comparison of AA cookies
after rejection vs after revocation
(c) Comparison of AA cookies
after acceptance vs after revocation
Figure 4: Change in the number of advertising and analytics (AA) cookies across different options
In this section, we address RQ2 and measure whether cookies requiring consent (Advertising and Analytics, or AA cookies) are deleted upon user’s consent revocation. As described in Section 4.1.2, we recorded the cookies of websites in four stages – upon initial landing, after accepting consent, after rejecting optional cookies, and after revocation. Out of the 120 websites where revocation was possible within the same interface, we find AA cookies after revocation on 69 (57.5%) websites.
Figure 4 compares the change in the number of AA cookies on websites after revocation w.r.t. other stages. Surprisingly, the number of AA cookies increases on most websites after revocation w.r.t. initial landing and w.r.t. after rejection (see red bar in Figures 4(a) and 4(b)). Additionally, on the majority of analysed websites, the number of AA cookies after acceptance remains the same after revocation (see pink bars in Figure 4(c)). Websites that add or keep AA cookies after consent revocation violate LR4 (Revoking requires stopping data processing and deletion of consent-based data), P1 and P2. Such AA cookies are processed unexpectedly and contrary to user’s decisions, without a legal basis and are thus illegal  [1, Art. 6(1)(a)].
7 Validity and Consistency of Consent
In this section, we analyse how consent information is stored and shared behind a website interface, whether it is consistently stored across different storage and APIs and whether it is legally valid, thus answering RQ3. We analysed 281 websites with detected CMPs (DS2 and DS3 from Table II) to check the validity and consistency of the consent string that these CMPs implement. Together with a legal expert co-author of the paper, we analyzed 11 purposes predefined within IAB Europe TCF v2.2 [35] (see Table IX in the Appendix) [^4] These purposes are largely consistent with those of the TCF v2.0 [37].. We determined that purposes 2-9 require consent and demand user action to be selected, thus such purposes cannot be enabled by default in TCString. Purposes 10-11 are exempted from this requirement and may be enabled by default. Purpose 1 is storage-based and enabled by default.
Positive consent: if the given TCString contains at least one of the purposes from 2-9, and at least one vendor in its vendor list, we consider it to contain a positive consent. Such consent is correct only within the Acceptance phase, where the user has actively selected purposes 2-9 requiring user action.
Negative consent: if only purposes 1, 10 or 11 are enabled in the TCString, we conclude that it contains negative consent because none of these purposes require any user action as per our legal analysis. Consequently, if such TCString is present upon initial landing, or after rejection or after revocation stage, we consider consent to be registered correctly.
For the consent strings extracted from OTAG variable, or from OptanonConsent cookie, we follow the OneTrust specification [70] and extract purpose numbers from its “groups” parameters. Since we do not have a precise specification of the meaning of these purposes, we cannot analyse which ones require consent. We therefore assume that a consent string contains a negative consent if it matches the value observed upon the initial landing stage, and a positive consent if the value contains more purposes than at initial landing.
Storage method or API Number of websites Positive consent
Total using
the method
Rejection
possible
Revocation
possible
Initial
landing
After
rejection
After
revocation
__tcfapi 163 150 136 1 2 17 (15 not updated)
TCF Cookies 130 104 101 0 0 12 (11 not updated)
TCF localStorage 41 31 32 0 1 8 (7 not updated)
OTAG 176 164 152 0 10 22 (16 not updated)
OptanonConsent Cookies 131 130 130 0 0 13 (10 not updated)
TABLE V: Websites where consent strings in browser storage and returned by APIs were positive in different phases.
7.1 Results
7.1.1 Positive consent upon initial landing, after rejection and after revocation
We first analyse whether the TCStrings returned by the __tcfapi can be considered as a “baseline” in our further comparison, that is whether such TCStrings are valid. For this purpose, we analysed whether the __tcfapi returns a positive consent at three stages: Initial landing, Rejection and Revocation, as shown in Table V.
After Rejection and Initial landing: Only two prior works [4, 40] analysed whether positive consent is stored or returned by APIs upon Initial landing and After Rejection. Compared to previous works, we found very few cases where positive consent is present after rejection in websites implementing the TCF (only 3 websites out of 150). In contrast, 10 websites, out of 164, store positive consent after rejection with OTAG.
After Revocation: We detected multiple websites where a positive consent is present after revocation, which has not been studied in previous works. In the websites with the TCF, 17 (12.5%) out of 136 websites provide positive consent in the __tcfapi, 12 (11.9%) out of 101 websites store a positive consent in a cookie, and 8 (25%) out of 32 websites store a positive consent in a localStorage. In the websites using OneTrust, we observe 22 (14.5%) out of 152 websites with positive consent returned by OTAG , and 13 (10%) out of 130 websites with positive consent in the OptanonConsent cookie. Positive consent constitutes potential violations of LR5 (Correct Consent Registration) which entails that the registered consent must be identical to the user’s choice in the user interface; and of the principle P2 since websites did not implement technical measures and safeguards to assure that revocation is done efficiently.
7.1.2 Consent strings are often not updated after revocation
Since a positive consent was present after revocation in numerous cases, we analysed whether consent strings were indeed properly updated when the user revoked consent on these websites (Table V). On websites implementing __tcfapi, we observe that 15 (9.3%) websites did not update the consent string, while 11 (10.9%) and 7 (21.9%) websites did not update the TCString stored in cookies, and in the localStorage respectively. Two websites that updated consent returned by __tcfapi are sourceforge.com and ft.com. Website sourceforge.com did not have a banner, stored a negative consent at initial landing, but updated it into a positive consent upon revocation. Website ft.com only updated the “consentScreen” parameter (which is unrelated to purposes) in the TCString after revocation, thus leaving a positive consent. Among websites implementing OTAG, 16 (10.5%) out of 152 did not update consent after user revocation. 6 websites have updated the consent string by adding or removing only some of the purposes, but the resulting consent string still contained a positive consent. Curiously, 3 out of 6 websites – cisco.com, opendns.com and webex.com – added one purpose to the consent string after revocation.
7.1.3 Inconsistency among consent strings between browser storage and APIs
Table VI shows the number of websites where the TCString found in the browser storage (cookies and localStorage) is not consistent with the TCString returned by the __tcfapi, and where the OptanonConsent cookie does not match the consent value returned by the OTAG. Such mismatches result in either an incorrect storage of consent or incorrect functioning of the APIs, and result in wrong consent being returned to third-parties. 5 websites implementing the TCF returned different TCStrings from __tcfapi and the TCString stored in the cookie after revocation. In 2 of such websites, (freep.com and megaphone.com) the TCString in the cookie was not updated, and still contained a positive consent, while the TCString returned by the __tcfapi was properly updated into a negative consent. In the website aol.com, while the __tcfapi returned a positive consent, the value in the cookie stores negative consent.
In 3 websites (reuters.com, manchestereveningnews.co.uk and portfolio.com), the TCString returned by the __tcfapi showed a negative consent, while the TCString stored in the localStorage remained unchanged, i.e, it contained positive consent. On websites implementing OneTrust CMP, we observed 4 mismatches between the OTAG value and the OptanonConsent cookie value, where the cookies stored a negative consent and were shared on the network. All these mismatches can lead to the incorrect consent being shared to third-parties.
These detected mismatches between consent strings across browser storage and APIs can be due to the fact that websites do not update, consistently, browser storage and APIs when users revoke consent. Consequently, websites do not correctly register consent, as they are obligated to, in order to assure that the registered consent is identical to the user’s choice in the user interface, thus, infringing the consent requirement LR5. Moreover, the principle P2 is not also complied with since websites did not implement technical measures to assure that revocation is done efficiently.
Browser storage vs. API # websites
TCF cookie vs. __tcfapi 5 (100)
TCF localStorage vs. __tcfapi 3 (32)
OT cookie vs. OTAG 4 (130)
TABLE VI: Inconsistent consent strings across browser storages and APIs after revocation. The numbers in brackets show the number websites that used both the storage and the API.
7.1.4 Inconsistency between consent returned via __tcfapi and consent shared on the network
We compare consent strings from __tcfapi with the consent strings found in outgoing network requests and incoming responses for two stages: Acceptance and Revocation. We consider the cases where the TCString returns the correct consent, i.e, a negative consent in case of revocation.
We investigate inconsistencies in the consent strings shared to and received from the third-parties, which, in turn, helps identify the responsible party for such inconsistencies. Previous works have only partially analysed such inconsistency [4, §VIII.A], where positive consent was found to be sent within a specific gdpr_consent URL parameter on websites where __tcfapi did not contain a positive consent string. We however observed that the consent can be shared via differently named URL parameter as well. Additionally, the consent string can be sent in the POST data as part of the request as well. We find different inconsistencies on 8 distinct websites out of the 136 websites where revocation is possible.
These inconsistencies could arise due to three reasons. (1) Due to delay in updating the consent after user revokes consent; (2) due to some scripts not being updated about revocation in consent; (3) due to introducing/using a different TCString. We do not consider delay in sending the updated consent as a violation since it is implementation specific. However, there are 8 websites where we observe inconsistencies causing possible violations. Out of these 8 websites, 4 of them (forbes.com, time.com, n-tv.de and cadenaser.com) had a different TCString on the network. In the remaining websites (deadline.com, kotaku.com, manchestereveningnews.co.uk and walesonline.co.uk), we observed that the old positive consent string was shared on the network by some scripts even after revocation. Details regarding these inconsistencies can be found in Appendix .1.
The detected mismatches between consent strings returned by __tcfapi and the TCString shared on the network (Table VIII) lead to an incorrect registration of consent revocation by websites and also by the implemented CMPs. Consequently, websites do not comply with LR5 since they do not correctly register user revocation, as they are obligated to, in order to assure that the registered consent is identical to the user’s choice in the user interface. As a result, a single TCString does not serve as a proof of revocation, which should consist of a negative consent. Moreover, the principle P2 is also not complied with since websites did not implement technical measures to assure that revocation is done efficiently.
8 Evaluating Correct Communication of Consent Revocation to Third-parties
In this section, we address RQ4 and investigate if all the third-parties that were informed of the consent acceptance (either by accessing APIs or via HTTP requests) are also informed of consent revocation.
8.1 Results
8.1.1 Not all third-parties are informed of revocation by accessing consent via APIs
To examine which third-parties access the APIs implemented by CMPs, we override the implementation of __tcfapi and OTAG (see §4.2.4). Overall, we found 23 (9.6%) websites out of 238 where at least one third-party accessed the API to fetch positive consent after acceptance, but did not access the API to fetch revoked consent.
Among these, on 163 websites, where __tcfapi is implemented (see Table V), on 14 (8.6%) websites at least one third-party requests the __tcfapi with the command “getTCData” after acceptance but not after revocation [^5] In the latest version of TCF, there is an option for third-parties to add an event-listener for consent update. However, since the implementation of this API is not standard, we could not intercept calls to it.. Out of 176 websites that support OTAG, on 13 (7.4%) websites, at least one third-party accesses the OTAG variable after accepting, but not after revocation. On 4 of these 13 websites, scripts from the third-party domain ads-static.conde.digital read the value of the variable, indicating that this variable can be used to know the consent choice given for AA purposes, since this domain is present in EasyList [72].
While CMPs offer APIs and other methods for third-parties to access consent information, third-parties become responsible for requesting access to consent and for being informed when consent is revoked. Therefore the correct implementation of proper event listeners by the website or by CMPs becomes necessary to inform third-parties about consent revocation, specially since consent can be updated multiple times in a single user session.
% 3rd-parties not informed Number of websites
<25absent25<25< 25% 1
[…]
100100100100% 45
TABLE VII: Percentage of third-parties informed of acceptance but not informed of the revocation in 101 websites.
8.1.2 Not all third-parties are informed of the consent revocation via HTTP requests
On each website where consent revocation is possible, we detected third-parties that are informed of consent information via the TCStrings in the URL or postData (see §4.2.4). We then detect third-parties informed of consent after acceptance but not informed of revocation.
Of the 136 websites that implement __tcfapi and support revocation, 101 (74.2%) websites contain at least one third-party that was informed of consent after acceptance, but not after revocation via HTTP requests. Moreover, on 68 of these websites, third-parties that were not informed about revocation set cookies through the “Set-Cookie” HTTP header, and thus, did not stop processing user’s data after revocation. Table VII shows the percentage of third-parties not informed of consent after revocation. Surprisingly, 45 (44.5%) websites did not inform any of the detected third-parties, while 35 (34.6%) of them did not inform more than 75% of included third-parties. Figure 5 shows most prevalent third-parties that were not informed of the revoked consent via HTTP requests but were informed of consent acceptance. These websites include the big ad-tech industry actors doubleclick.net, criteo.com and adnxs.com. Interestingly, 73.77% of all third-party domains that do not receive a communication about consent revocation are present in EasyList [72] and therefore participate in advertising or tracking.
Figure 5: Top 15 third-parties not informed of the revoked consent via HTTP requests and the number of websites including these scripts but not informing them of revocation.
8.1.3 Failure to communicate consent revocation results in unlawful data processing
Our findings indicate that third-parties are not informed about consent revocation. This fact entails that these tracking- and advertising-based third-parties may continue to unlawfully process user’s personal data even after users have revoked consent on the website’s interface, since they were not updated of the user’s decision by the website. Consequently, websites from both cases (communication of the TCString via API and HTTP requests) are in potential violation with LR6 (Communication of withdrawal to third-parties) that demands websites to communicate consent revocation to the third-parties. The continuous unexpected processing of user’s data after revocation infringes the fairness principle P1 and the data protection by design principle P2 due to the lack of technical measures and safeguards efficiently enabling the exercise of the revocation right by websites.
9 Recommendations
In this section we propose recommendations to regulators for enhancing the implementation of consent revocation and ensuring legal compliance, in the light of our findings.
R1. Need to unify EU requirements for consent revocation interface. Our research shows that websites implement revocation inconsistently across the Web (see Table IV). For example, the text placed in the website footer differs across websites (e.g., “Cookie Settings”, “Privacy Settings”, but also “EU Privacy”), and in many cases does not mention revocation. It is even harder to locate revocation links within privacy policies that are labelled with text such as “How do I control cookies and how my data is used?”, or “Managing our analytics cookies”. The vague and ambiguous representation of these options misleads the users and prevents them from exercising the revocation right which should be clearly and distinctly recognisable (see LR1). Icons, that are supported by regulators and are easy to locate, use different visualisations, which may also confuse the user. EU DPAs should propose unified interface requirements describing the interface, location, and wordings of revocation. An inspiration can be taken from CPRA regulation in California, where three acceptable wordings are proposed, which simplifies locating it with automated means to measure compliance [41].
R2. Need to standardize consent storage. We observed various inconsistencies in storing the TCStrings and OneTrust consent strings (see Section 7). This is partly due to the usage of different storage options that are not updated simultaneously when consent is revoked. We invite regulators to establish standards for consent storage, and implement security measures to protect the integrity of consent strings.
R3. Need to standardize consent communication via event listeners. Our results show that only 43 out of 163 websites had third-parties installing event listeners in order to be updated about the change in consent decisions (see Section 8.1.1). While OneTrust suggests developers to add event listeners to synchronize consent updates when integrating advertising systems [73], our results show that IAB Europe TCF does not standardize event listeners, prompting every CMP to propose their own solution. This makes it harder for third-parties to adapt to different implementations across websites, placing the burden on them along with legal consequences (for continuous processing of personal data), without offering a practical solution to ensure compliance. We therefore recommend standard-setting bodies like IAB Europe TCF to standardize the implementation of event listeners and the callback functions returned by event listeners to inform third-parties about consent revocation. We also propose regulators to take a position on the means of communicating consent to third-parties. We believe that event listeners are able to provide reasonable means for communicating revocation decisions, and help in the allocation of responsibility to collect valid consent by third-parties present on a website.
R4. Need to regulate consent communication via HTTP requests. The most prevalent violation we detected indicates that 74.2% of websites do not properly inform all third-parties about consent revocation when the HTTP request method is used (see Section 8.1.1). Our result shows that this concrete way of informing third-parties exempts big advertising actors, such as doubleclick.net and criteo.com (see Figure 5) from any responsibility regarding data deletion when users revoke consent. Moreover, these third-parties only receive positive consent from the majority of websites and consequently, while illegally processing data from users who revoked consent, these companies can falsely demonstrate evidence of compliance to regulators by providing a false proof of consent. We therefore propose that EU regulators express their position regarding how websites should inform third-parties, and whether HTTP requests is an acceptable method, since our findings demonstrate that this approach leads to the most prevalent violations.
10 Data Availability & Responsible Disclosure
Data availability. As described in Section 4.1 and Figure 1, we collect background data, screenshots and label the websites during data collection. Screenshots and labels help in categorize the websites into different categories, "via same interface", "via different interface" and "no revocation". A few examples can be found in the Appendix, in Figures 6 and 7. A comprehensive list of these websites is available in the supplementary material containing the dataset and crawler [74]. The network logs to identify inconsistencies and code of the crawler to collect the data are also available in the supplementary material.
Responsible disclosure. As of the date of submission of this paper, 14 November 2024, we have notified all the companies who own the domains that we explicitly mention in the main text of our paper. These domains were mentioned in Sections 57 and 8.1.1. We will add information on the responses we will receive upon acceptance of this paper.
11 Conclusion
In this work, we propose a framework to audit compliance of consent revocation on the Web. We found multiple instances of violations of the EU Data Protection law while analysing interfaces of revocation, including use of different interfaces (19.87%) and more effort to revoke than to accept (20.5%). Multiple violations were observed on the usage of cookies and storing positive consent despite user’s revocation. Most shockingly, on 74% of websites, third-parties that have received consent upon user’s acceptance, are not informed of user’s revocation, leading to the illegal processing of users’ data by such third-parties. Our findings emphasise the need for improved legal compliance of consent revocation, in particular, proper, consistent, and uniform implementation of revocation communication and data deletion practices.