### v3

# Johnny Can’t Revoke Consent Either: Measuring Compliance of Consent Revocation on the Web

Abstract.
The EU General Data Protection Regulation (GDPR) requires websites to facilitate the right to revoke consent from Web users. Prior works have examined consent management by auditing that user choices are correctly stored, and comparing cookies set upon acceptance versus rejection to assess compliance. While these studies measured compliance of consent with respect to the various consent requirements, no prior work has studied consent revocation on the Web. Therefore, it is unclear how difficult it is to revoke consent on the websites’ interfaces, and whether the revoked consent is properly stored and communicated behind the user interface.
Our work aims to fill this gap by measuring compliance of consent revocation on the Web on Tranco’s top-200 websites. We found that 19.87% of websites make it difficult for users to revoke consent throughout different interfaces, 20.5% of websites require more effort than acceptance, and 2.48% do not provide consent revocation at all, thus violating EU legal requirements for valid consent. 57.5% websites do not delete the cookies after consent revocation enabling continuous illegal processing of users’ data.
Further, we analyzed 281 websites implementing the IAB Europe Transparency & Consent Framework, and found 22 websites that store a positive consent despite user’s revocation. Surprisingly, we found that on 101 websites, third parties that have received consent upon user’s acceptance, are not informed of revocation, leading to the illegal processing of users’ data by such third parties according to EU laws. Our findings emphasize the need for improved legal compliance of consent revocation, and proper, consistent, and uniform implementation of revocation communication to third-parties.
journalyear: 2025journalvolume: YYYYjournalnumber: Xdoi: XXXXXXX.XXXXXXX
1. Introduction
In recent years, compliance with privacy and data protection laws on the Web has gained a lot of attention, both from the research community and regulators. The EU Data Protection framework, consisting of the General Data Protection Regulation (GDPR) (European Parliament and Council of the European Union, 2016) and ePrivacy Directive (ePD) (ePD-09, 2009) sets requirements for a legally-valid consent when tracking technologies are deployed on a website or app. Such consent is usually implemented via consent banners and must satisfy seven requirements for validity: it must be prior to any data collection, freely given, specific, informed, unambiguous, readable and accessible, and revocable (European Parliament and Council of the European Union, 2016, Art. 4(11), 7)(Santos et al., 2020, para 17). While numerous studies have evaluated compliance of consent banners with the various consent requirements, such as the presence and visibility of the rejection button  (Matte et al., 2020a; Bollinger et al., 2022; Nouwens et al., 2020), one aspect has so far been overlooked by the research community and the regulators: the requirement of revocable consent, also known as consent withdrawal, which entails that the user has the option to change a prior preference regarding trackers due to the reversible nature of consent decisions (Directorate General Justice, European Commission, 2013).
According to the GDPR, users have the right to revoke consent at any time (European Parliament and Council of the European Union, 2016, Art. 7(3), Rec. 42). Accordingly, websites must facilitate the exercise of this right (European Parliament and Council of the European Union, 2016, Art. 12(2)) by providing an option to revoke consent. Upon revoking consent, websites must subsequently comply with an additional obligation to delete data previously processed on the basis of that consent, and without undue delay even if the user did not explicitly exercise their right to request data deletion (European Parliament and Council of the European Union, 2016, Art. 17(1)(b)). This paper focuses exclusively on the right to revoke consent. Websites that do not allow consent revocation are deemed to be processing data illegally, and run the risk of being fined for not complying with the legal requirements (NOYB, 2024; for Digital Government, 2023; CNIL, 2023; (2019), UODO; DPA, 2024).
However, it remains unclear whether websites provide users with compliant methods to revoke their consent, whether revoked consent is properly stored by the websites behind the consent interface, and whether it is communicated to third-parties that collected users’ data. Our work aims to fill this gap by measuring the compliance of consent revocation on the Web, addressing the research questions:
  • RQ1:
    Are the revocation interfaces on websites compliant with EU data protection laws?
  • RQ2:
    Do advertising and analytics (AA) cookies that require consent get deleted once consent is revoked?
  • RQ3:
    Is consent stored correctly in the browser, and is it consistent across browser storage and APIs implemented by Consent Management Platforms (CMPs)?
  • RQ4:
    Are all third parties that are initially informed of “acceptance” also notified when the consent is revoked?
To address these questions, we have set up a team of several computer scientists and a legal expert to measure compliance of consent revocation on the Web, making the following contributions:
  1. (1)
    We provide the first in-depth legal analysis of the GDPR, ePD and other legal guidelines for consent revocation interface, storage and communication, and establish six operational legal requirements (§3).
  2. (2)
    We provide a method to audit compliance of consent revocation interfaces (§4.1), and apply our analysis on Tranco’s top-200 websites identifying multiple instances of non-compliance with the legal requirements (§5).
  3. (3)
    We study the effect of revoking advertising and analytics (AA) cookies, finding that on the majority of websites, revocation does not decrease the number of AA cookies, thus violating the GDPR (§6).
  4. (4)
    We propose a methodology (§4.2) to evaluate revocation on websites that implement IAB Europe Transparency and Consent Framework (TCF) and the most popular Consent Management Platform, OneTrust. We analyse 281 such websites and evaluate the storage of consent behind the consent banner interface (§7) and the communication of revocation (§8), thereby detecting multiple non-compliant practices.
  5. (5)
    Finally, we give recommendations to further improve consent revocation compliance (§9).
Violations within the interface and cookies in the top-200 websites Prevalence
Revocation via a different interface/medium (§5) 20.25% (32/158)
Two or more steps to find revocation interface vs zero steps to accept (§5) 20.8% (33/158)
Data processing based on AA cookies is not stopped upon user’s consent revocation(§6) 57.5% (69/120)
Violations beyond the interface in 281 websites with CMPs Prevalence
Positive consent after revocation (TCF-based consent string) (§7) 16.17% (22/136)
Positive consent after revocation (OneTrust-specific consent string) (§7) 14.47% (22/152)
Third-parties informed of consent acceptance via HTTP requests, but not informed of revocation (§8) 74.2% (101/136)
Table 1. Summary of most prominent potential violations of consent revocation implementation. In brackets, we mark the number of websites with violations and the total number of websites analysed for each case.
Table 1 summarizes the most prevalent (more than 10%) potential violations of the GDPR and ePD that we detected while analysing both the consent revocation interface, cookies, and the consent storage and sharing with third-parties. Among the top-200 websites, we find that 22.7% (36) websites are compliant with GDPR and ePD within their revocation interfaces and management of cookies (see remaining potential violations in rows 1-3), and the compliance is more common in websites that employ CMPs. We, therefore, further investigate websites with CMPs for possible non-compliance. Of the 281 with CMPs, 251 provided revocation options – among them, we further check if a positive consent is stored even after revocation and if the consent modification is communicated to third-parties that were informed of acceptance. Overall, we find that 52.6% out of 251 websites providing revocation do not have a possible violation beyond the interface, while the remaining websites either had positive consent after revocation or communicated an initial positive consent to third-parties but did not communicate negative consent after revocation (rows 4-6).
2. Background and Related Works
In this section, we provide background on different techniques for consent management and discuss prior works on compliance in the context of consent banners, opt-out, and revocation.
Consent Interfaces and Compliance. Consent banners have become a common method for obtaining consent for online tracking. Recent works (Santos et al., 2020; Bielova et al., 2024; Degeling et al., 2019; Hils et al., 2020; Kampanos and Shahandashti, 2021; Trevisan et al., 2019) have examined consent banners following the implementation of the GDPR and the ePD while numerous others (Gray et al., 2021; Giese and Stabauer, 2022; Ma and Birrell, 2022; Soe et al., 2020; Mathur et al., 2021; Machuletz and Boehme, 2020; Sanchez-Rola et al., 2019; Matte et al., 2020a; Nouwens et al., 2020; Leenes and Kosta, 2015; Aziz and Wilson, 2024) explored the impact of privacy regulations on consent design, and how the websites and CMPs might violate these laws, e.g., by using dark patterns. Tools have been proposed to help users handle consent banners automatically (I dont care about cookies, [n. d.]; Ninja Cookie, [n. d.]; AU, 2022; superagent, [n. d.]; Bollinger et al., 2022; Jesus and Pandit, 2022; Kulyk et al., 2018; Pandit et al., 2019). Several works (Bollinger et al., 2022; Hils et al., 2020; Nouwens et al., 2020; Kirkman et al., 2023) focused on automated analysis of consent banners at scale.
IAB Europe TCF and Compliance. The current “de-facto” standard for the consent banners in the EU is the IAB Europe Transparency and Consent Framework (TCF) (Europe, 2025). The current TCF v2.2 (IAB, 2025b) defines the pre-defined purposes for processing personal data by Consent Management Platforms (CMPs). If also defines the format to store the user’s choice, called TCString,that includes:(a) list of enabled third-party vendors registered within the TCF; (b) list of enabled purposes among the pre-defined purposes (IAB, 2025a).
The TCF standard does not specify which of the purposes require user consent. According to GDPR and ePD, only purposes that do not require user consent (and users’ explicit interaction with the CMP) may be enabled in TCString  by default. Matte et al. (Matte et al., 2020b) analysed which purposes in TCF v1.1 and v2.0 require consent according to the GDPR and ePD; however, this analysis was not extended to TCF v2.2, which we perform in this paper.
Once a user has made their choice on the CMP interface, the CMP is required to store the corresponding TCString in the browser (however, the specific storage mechanism is not specified in TCF v2.2), or implement an API for third-parties to check the TCString. The TCString can be accessed via the __tcfapi function by third-parties in first-party context or via the JavaScript postMessage API to communicate with a special __tcfLocator iFrame. Following TCF v1.1 (Lab and Europe, 2018), TCString can also be shared in the outgoing HTTP requests via URL-based methods. However, TCF v2.2 does not specify any URL parameters to be used for this.
Prior works (Matte et al., 2020a, b; Smith et al., 2024; Zhang et al., 2024) have focused on the analysis of IAB Europe TCF and its consent implementation. However, none of these works studied how CMPs manage consent revocation, which is one of the main objectives of our work.
Consent Rejection and Revocation. While multiple studies (Matte et al., 2020a; Nouwens et al., 2020; Bollinger et al., 2022) detect and analyze the presence of a reject button on consent banners, none of the prior works evaluated the interfaces that allow revocation of a given consent within the EU data protection framework. Nevertheless, since the updates of the CPRA regulation in California, multiple studies have analysed the implementation of the right to opt out of selling or sharing user data within websites. Tran et al. (Tran et al., 2024) propose a methodology to measure compliance of opt-out links’ wordings automatically found on websites subject to CCPA and CPRA (California State Legislature, 2018). Liu et al. (Zengrui Liu and Saxena, 2024) measure the impact of opting out in the presence of advertisers and tracking mechanisms. Aziz et al. (Aziz and Wilson, 2024) assessed the IAB CCPA Compliance Framework (analogous to IAB TCF in Europe) and detected that opt-out signals are not honored on websites. A similar latest work by Du et al. (Du et al., 2024) detected violations regarding withdrawal interfaces in mobile applications.
Other studies evaluated user perceptions when interacting with opt-out mechanisms. As such, Habib et al. (Habib et al., 2019) evaluated the usability of data deletion and opt-out options related to email communication and targeted advertising. Habib et al.(Habib et al., 2022) evaluated how different instructions about revocation in the banner’s text impact users’ ability to find revocation options.
None of these works, however, studied whether the revocation interfaces comply with the EU laws. Additionally, no prior work has shed light on whether websites correctly store and manage revoked consent, and inform all the third-parties that were previously informed of consent acceptance.
3. Legal Compliance for Revocation
The EU Data Protection Framework provides legal principles and requirements that can be applied to websites and specifically to consent revocation. The GDPR applies to the processing of personal data (Board, 2007), and the ePrivacy Directive (ePD) (ePD-09, 2009) provides supplementary rules, particularly, for the use of tracking technologies ((2019), EDPB). Whenever tracking data is stored and read from the user’s device, the ePD (ePD-09, 2009, Art. 5(3)) requires websites to request user’s consent when tracking is used for purposes not strictly necessary for the service provided, e.g., advertising (European Data Protection Board , Article 29 Working Party(2012), EDPB; (2013), EDPB); purposes required for a website to operate are exempted (ePD-09, 2009, Recital 66). The way to assess with certainty whether consent is required is to analyze the purpose of each tracker on a given website ((2013), EDPB; Fouad et al., 2020).
The GDPR, as well as guidelines from the Data Protection Authorities (DPAs) and from the European Data Protection Board (EU advisory board, representing all EU DPAs) provide legal requirements for consent revocation. While guidelines are not legally-binding, they are part of the EU framework for data protection which we apply in this work to discern when revocation methods are compliant. A website can be held liable and fined if it fails to comply with the GDPR principles and requirements for valid consent, including revocation. Relevant GDPR principles are presented below and numbered with P, and legal requirements for consent revocation for websites are numbered with LR.
P1 (Fairness).
Websites must not process personal data in a unjustifiably detrimental, discriminatory, unexpected or misleading way (European Parliament and Council of the European Union, 2016, Art.5(1)(a))(European Data Protection Board, 2020a, para 17).
P2 (Data protection by design).
Websites must implement organisational measures and safeguards efficiently to enable the exercise of the revocation right (European Data Protection Board, 2020b, para 68), (European Parliament and Council of the European Union, 2016, Art. 25(1)); (European Data Protection Board, 2020b, para 70). Revocation options should be provided in an objective and neutral way, avoiding any deceptive or manipulative language or design (European Data Protection Board, 2024, 2022, para 16). If a website requires more effort to revoke than to give consent, it is deploying a dark pattern (European Data Protection Board, 2022, para 30).
P3 (Accountability).
Websites must demonstrate that revocation is performed easily and effectively (European Parliament and Council of the European Union, 2016, Arts. 5(2), 24(1), Rec. 74).
4. Methodology
Next, we describe the data collection and analysis methodology to answer our research questions. We built a semi-automated crawler based on a Selenium-instrumented Chromium v122.0.6261.128. We used the crawler to collect data between March and June 2024, within the EU. Table 2 presents an overview of the three datasets we collected, which are further explained in this section.
Figure 1. Data collection pipeline: to address RQ1, we collect screenshots and label a website as shown in violet boxes; for RQ2 we collect cookies at each stage and also at Rejection stage in a similar way (not shown in this figure).
4.1. Datasets for RQ1 and RQ2
Domains Collected dataset RQs
200
(ranked 1-200)
DS1: 105 domains without
CMPs
RQ1, RQ2
DS2: 56 domains with
CMPs
RQ1, RQ2,
RQ3, RQ4
1000 random
(ranked 201-5k)
DS3: 225 domains with
CMPs
RQ3, RQ4
Table 2. Datasets used for RQs.
We have implemented the data collection pipeline shown in Figure 1 to capture the elements of the user interface that need to be interacted with in order to exercise consent revocation. At each stage of our analysis, we save background data to further analyse cookies, consent storage, and communication. Figure 2 presents the details of the background data collection.
4.1.1. Website selection
We analyzed the revocation interface options on Tranco top-200 domains (Le Pochat et al., 2019; Tranco, 2023). 158 domains were reachable while 42 did not display a webpage because they were either CDNs, blocked or returned failures. We detected CMPs by decoding the TCString returned by __tcfapi following the methodology of Matte et al. (Matte et al., 2020a), after accepting consent. We detect the websites using OneTrust CMP by the presence of OptanonConsent cookie, specific for this CMP. Of the 158 domains, 56 used TCF and/or OneTrust (labeled “with CMP”), which we could detect.
4.1.2. Analyzing interfaces for consent revocation (RQ1)
We collected the data (including cookies, localstorage, request-response pairs and screenshots) in a semi-automated manner as shown in Figure 1.Manual effort was required to browse and locate revocation options, while all technical data (cookies/storage, screenshots, logs) were collected automatically. Automating this would require robust ML/NLP models, which may not generalize well due to the diversity in UIs and revocation paths, as discussed in results of RQ1, section 5.1.1. While we manually visited and navigated the websites, the data was collected automatically in the background. We took screenshots to record the user interfaces involved in reaching a revocation option. While recent works (Khandelwal et al., 2023, 2021; Bouhoula et al., 2024) provide automated tools to select initial consent options on websites (specifically, button and toggle HTML elements) based on user choices, our analysis shows that revoking consent may require navigating multiple links on the page to reset the options once the initial banner is closed. Moreover, the location of the consent modification options on websites are non-standard, and it is not clear whether (and where) such options are present. Therefore, revocation options are difficult to find and interact with automatically due to the heterogeneity of revocation implementations. We collected this data in four stages:
(1) Initial landing: We visit each website afresh with storage cleared for all websites in the browser, and collect the screenshots showing the presence of a banner.
(2) After acceptance: We record whether a banner is present and if an option is provided to the user to accept consent, in which case we accept all cookies.
(3) After revocation: After accepting consent, we check if an icon for modifying consent is visible on the page. If not, we check if an option is available in the footer or elsewhere on the main page. Finally, we check if the option is available on a different page reachable from the main page. For each of the options, we record the number of steps needed for revocation. If no option is found, we record that revocation is not possible.
(4) After rejection: To compare the behavior of the website after initial rejection (opt-out) with that after the consent is explicitly revoked, we clean the browser (includes clearing history, cookies, cache, passwords, form data, and site settings to ensure consistency in each measurement) and record the options to reject consent on the website. If a consent banner is present, we reject all cookies.
Amid all these steps, we also take screenshots in an automated manner at different stages by scrolling to the bottom of the pages or navigating to the settings in order to get the important information like keywords leading to revocation. Using these inputs and screenshots, we categorize the websites based on the revocation options, where such options are provided, and the number of steps required to revoke consent. We then map these categories to the legal requirements for GDPR and consent revocation (from §3) to identify and measure potential violations of the law. To reduce manual bias in our analysis, labeling was conducted by one author (twice, three months apart, in March and June of 2024.), with category examples co-developed with a legal scholar. Several examples were jointly reviewed by both a legal expert and a computer scientist.
4.1.3. Effect of revocation on AA cookies (RQ2)
In the background, we collect cookies stored in the browser at each of the four stages. Since only some cookies require consent (see §3), we followed Bouhoula et al. (Bouhoula et al., 2024) and consider that only cookies classified as Advertising or Analytics (AA) by CookieBlock (Bollinger et al., 2022) require user consent. If we detect AA cookies upon initial landing, after rejection, or after revocation, we map it to legal requirements and identify a potential violation of the law. Similarly, we analyse whether the number of AA cookies increases after revocation w.r.t. other stages of our analysis.
In this analysis, we do not include the websites without a banner because we need to compare cookies after explicit acceptance and after revocation. Websites without a banner do not allow a user to accept consent, and therefore, modifying consent on such websites would not mean “revocation” in legal terms.
4.2. Datasets for RQ3 and RQ4
In websites that allowed revocation, different methods were used to store the modified consent and communicate it over the network. To detect the presence of CMPs that implement the IAB TCF (Europe, 2025), and therefore use a standardised format for storing consent (Transparency and Consent String or TCString), prior works (Matte et al., 2020a; Zhang et al., 2024) used the __tcfapi function provided by TCF. We, therefore, performed an automated preliminary study to test this detection method. We also identify the detection methods for the use of OneTrust, which is the most popular CMP on Tranco top-20k websites (Hils et al., 2021).
4.2.1. Preliminary study
We performed an automated crawl querying for __tcfapi, a function that must be provided to use TCF’s functionalities, on the top-200 domains (Tranco, 2023). Out of 158 websites that were reachable, we found __tcfapi on only 32 (20%) of websites [^1] Similar to prior work (Matte et al., 2020a), we found that __tcfLocator is present on only 1 website where __tcfapi is not present. We therefore do not test for this method here.. We therefore further explored other ways to detect CMPs by analysing the implementations of the most popular CMP (Hils et al., 2021), OneTrust. By examining websites with OneTrust and its documentation (OneTrust, [n. d.]), we found that this CMP (1) stores user consent in a very specific format in a cookie named OptanonConsent, and (2) maintains a JavaScript variable, OneTrustActiveGroups (OTAG), which can be queried to get the current consent string. However, the format of the consent string accessible through these methods does not respect the official IAB Europe TCF TCString format.
4.2.2. Website selection
We use the methods developed in the preliminary study to detect websites that use TCF and OneTrust from top-200 websites, resulting in dataset DS2 with only 56 websites with CMPs. We randomly chose an additional 1000 websites between the rank 200 and 5000 and detected 225 websites with either TCF or OneTrust resulting in a set of 281 websites (DS2 and DS3).
Figure 2. Data collection pipeline: to address RQ3 (validity and consistency of consent), the collected sets of consent strings are highlighted in orange, while for RQ4 (access and communication of consent to 3rd-parties), the sets of scripts parties involved in the communication of consent are highlighted in violet. The objects in green represent our methods for collecting the data, and the boxes in yellow are further described in Table 3.
4.2.3. Data collection
Figure 2 shows the data collection pipeline for RQ3 and RQ4. The data is collected during all four stages of the experiment (highlighted as “Save background data” in Figure 1). We collect the consent strings (both TCString and consent variable specific to OneTrust) for further analysis of validity and consistency of consent storage (RQ3), and complement it with the information about scripts that access consent strings via APIs, scripts that initiate the requests and third-parties with whom consent is shared for the analysis of communication of consent (RQ4).
4.2.4. Collecting stored consent strings and third-parties that access them
The current TCF v2.2, does not define the specific location where the user consent should be stored. Therefore, we collect all cookies and localStorage objects’ values that match the TCString format. Additionally, we collect the OptanonConsent cookie that OneTrust uses to store a user’s consent. We also inject scripts in the webpage to query __tcfapi and OTAG to analyse consent strings that third-parties would receive if they query these APIs and examine network logs to detect and store TCStrings shared in the requests and responses. All the collected consent strings are visualised in Figure 2 in orange color.
Third-parties can access consent strings via APIs or receive them via network requests. We capture (1) third-party scripts that fetch consent string via __tcfapi or OTAG; (2) HTTP requests that contain consent strings in TCString format or OTAG. However, some requests might not have been recorded since they do not have a well-defined format. We store all initiators and receivers of consent strings (see violet stores in Figure 2).
4.2.5. Detecting and extracting TCStrings and specific categories from OneTrust in network logs
To detect the TCString in network logs, we look at different parts of the request/response pair. Previous works have extracted TCStrings only from specific URL parameters , such as gdpr_consent (Matte et al., 2020a) following TCF v1.1. A follow-up work (Zhang et al., 2024) checked for outgoing TCString in all URL parameters of HTTP requests and also in the HTTP cookie headers since TCF v2.2 does not specify URL parameters to be used.
When parsing URLs, we extract all query parameters, for example for a URL https://www.site.com /zzz.jpg?ISBN=XXX&UID=ABC123, we first extract the values XXX and ABC123 and then check whether they correspond to the TCString format, according to TCF specification (IAB, 2024). The orange boxes in Figure 2 represent the functions we used to extract TCStrings from different parts of the network logs that are further explained in Table 3.
Since TCStrings do not have any protection mechanisms, any third-party that receives a TCString, can potentially modify it and return a different TCString back to the browser. To detect inconsistencies in TCString between requests and responses, we analyse postData in POST requests, and the response data (HTML and JSON formats) from the servers. In each case, we record the request initiator to identify scripts that sent the wrong consent strings. To analyse HTTP GET requests, we search for TCString in the URL parameters, while for POST methods, we extract the TCString from JSON object if the data is sent in this format. To analyse HTTP responses and detect when TCString is modified by the server, we analyse redirected URLs: if response type is JSON, we parse it; otherwise, we parse the returned HTML file to extract TCString from the script, image and iFrame elements.
Our approach uncovers more instances of TCStrings as compared to previous works (Matte et al., 2020a; Zhang et al., 2024) — we detect 4,445 TCStrings in URL parameters, 3,423 TCStrings in the postData, 52 TCStrings in JSON objects and 170 TCStrings in HTML files. We also record requests where the cookie containing TCStrings is sent to third-parties.
Func. in Fig. 2 Extracting TCString
parseURL (URL) Extract all query parameters from URL and keep all values that match a TCString format or contain OneTrust cookie categories.
parseJSON (obj) Extract all {key:value} pairs present in obj and keep all values that match a TCString format or contain OneTrust cookie categories.
parseHTML (file) Extract URLs for img, iframe and script in file and parse them to extract TCStrings or OneTrust cookie categories.
Table 3. Extracting TCStrings and OneTrust consent string
4.2.6. OneTrust-specific encoding of consent
Differently from other CMPs, OneTrust CMP does not use the TCString but uses a specific format to store and communicate consent: the consent can be stored as key-value pairs in two locations, either in the OptanonConsent cookie or in the OTAG variable. The cookie values contain encoded categories of cookies (e.g., “analytics” or “advertisement”), that follow the format such as “X:1;Y:0;Z:1”, where X, Y and Z indicate cookie categories and 0 or 1 indicate rejection or acceptance to use the specific category in the consent. The values of the OTAG list all the allowed cookie categories (called “active groups” in OneTrust documentation); however, the encoded format of such categories is not specified in OneTrust documentation. Our analysis of inconsistency between the OptanonConsent cookie value and OTAG variable (section 7) revealed that when two values are not consistent, OptanonConsent cookie value contained correct record of user consent. We therefore opted for extracting the allowed categories from the OptanonConsent cookies and search all such encoding (only X and Z in the example above) in the URL query parameters to record the consent strings being sent to third-parties.
However, the variable in the query parameter having the consent string is not consistent. This leads to false positives where the active groups are represented using numbers like ‘1’, ‘2’, and ‘3’ or “1:1,2:0,3:0,4:1". These representations are hard to distinguish from query parameters used for purposes other than sharing consent. We manually analyzed such websites, and hence, could have missed some websites due to human error.
4.2.7. Wrappers for __tcfapi and OTAG
To identify __tcfapi access, we override the function to record all the scripts that access it. We also record what command was requested, e.g., “ping”, “getTCData” or “getEventHandler”. A similar approach was adopted by Matte et al. (Matte et al., 2020a). However, they observed calls made to the APIs in an older version of TCF, and their analysis did not include OneTrust. We override getter and setter methods of the OTAG variable to record which domain is writing/updating it (mostly by a designated CMP script) and reading/accessing it.
4.2.8. Communicating consent modification
We check the network logs to investigate if third-parties that received the TCStringand One Trust consent string (containing allowed cookie categories) after acceptance had also received the updated consent after revocation. We do this by observing requests to third-party URLs and if the request contained the TCString or One Trust consent string either in the URL or in the POST data as described in Figure 2.
4.2.9. Identifying responsible parties
We record the initiators of all the network requests and use this information to identify which third-party script changed the consent. We match the script’s domain of provenance with the CMP name decoded from the TCStrings. If the name and domain do not match, in case of inconsistencies, we manually check the domains for cases where the CMP names are either shortened or have words like ‘privacy center’ or ‘CDN’. If we do not identify any such indicator that the third-party could be the CMP, we classify the responsible party just as a third-party.
4.3. Limitations
4.3.1. Accuracy of AA cookie classification
Even though CookieBlock (Bollinger et al., 2022) has better accuracy than other tools for classifying cookie purposes, it might not be entirely accurate: a website may declare some of the cookies, classified as AA, as necessary in their cookie policies. Such cookies, classified by CookieBlock, will be, nevertheless, labeled as AA and considered to require consent.
4.3.2. Not capturing network logs while searching for revocation option on secondary pages
When reaching the revocation settings takes multiple steps, for a few seconds, log-entries with the old consent (after acceptance) are also saved along with the network logs meant for “after revocation” stage. We remove these log-entries by searching for inconsistencies after the consent is revoked. However, the set of log-entries misclassified as “after revocation” are not checked for inconsistencies, and we may have missed a few cases similar to the ones mentioned in Table 8 in the Appendix.
4.3.3. Unable to capture access to consent when the third-parties use event listeners
Callback functions returned as an event listener by __tcfapi do not have a standard format or name to track the access requests made to the API. Hence, we do not track API accesses when event listeners are used.
4.3.4. Determining description of purposes in OneTrust
We do not know how the purposes displayed by OneTrust or any CMP used by the website are mapped to the user’s actual consent. Since there is no particular format used to store the consent string, it can vary between different CMPs and within websites using the same CMP.
4.3.5. Delay in registering user’s consent
While collecting network logs to assess the communication of consent revocation to third parties, we observed (see Section A) that some websites may take time to register that a user revoked their initial consent. To account for this, we exclude cases with delays in updating consent strings from our list of violations. However, we include cases where some third parties correctly include the updated consent string in their network requests, while others still send the old consent (positive consent string). As the third-party notification happens server-side, this may result in some false-positives on our end.
4.4. Data Availability and Disclosure
As described in §4.1 and Figure 1, we collect background data, screenshots, and website labels. The complete dataset, including the network logs and the crawler, is provided in the supplementary material (Anonymous, 2025). Some examples are shown in the Appendix.
As of the date of submission of this paper, 28 February 2025, we have notified all the companies who own the domains that we explicitly mention in the main text of our paper. We will add information on the responses we receive, in the final draft.
5. Revocation Interface and its Compliance
To address RQ1, we examined user interfaces for revoking consent and checked if they (potentially) violated EU legal requirements. We first categorize the $158158$ reachable websites (datasets DS1 and DS2 from Table 2), based on their consent banners: 108 (67%) websites display a consent banner to the users on their first visit, 8 (5%) websites displayed a banner with no option, not allowing to accept or reject consent, and 45 (28%) websites did not display any banner. [^2] This can be explained by the fact that these websites were hosted outside of the EU and did not yet implement compliance solutions with EU laws. We manually classify the 158 websites based on the interface to revoke consent: 120 (74.5%) websites provide users option to revoke consent within the same interface where the consent request took place, or navigates to related pages for revoking consent; 32 (19.8%) websites provided options via different interface, where such options are present outside the interface or medium, where the consent request took place, and 9 (5.6%) websites offer no revocation option.
Within the Same Interface Via Different Interface No Revocation
Banner #
Icon
Footer
Options
Banner
on Policy
Options
via Policy
Settings
or links
After
Login
Contact/
Email
No option
provided
Consent banner 105 8 61 4 23 8 1 0 0
No option banner 8 0 0 0 5 3 0 0 0
Without banner 45 0 7 2 7 17 0 3 4 [+5]
Total 158 8 68 6 35 28 1 3 9
Table 4. Prevalence and type of consent revocation options on 158 websites (rank 1-200) from DS1. We use different colors to represent the level of compliance of the detected practices. Green color represents zero or one steps, yellow represents two or more steps, and pink and red represent violations of varying severity.
Table 4 summarizes the results for these three categories. To further evaluate legal compliance, for websites that provide revocation within the same interface, we count the number of steps required to revoke consent based on Figure 1. Figure 6 (in the Appendix) shows example screenshots for each of these categories.
5.1. Results
5.1.1. Compliant revocation interface (zero or one steps)
Out of the 158 websites, only 8 (5.6%) offer a persistent icon or button floating on the page, thus requiring zero steps to reach the revocation option. $6868$ (41.6%) websites offered a link option in the footer of the page, requiring one step to reach the revocation option. 6 (6.8%) websites showed a consent banner or icon when accessing the privacy policy page from the footer, therefore also requiring one step to revoke consent. All these implementations found on 82 (51%) websites (labeled as Icon, Footer Options and Banner on Policy in Table 4) comply with revocation GDPR principles and consent requirements (LR1-3) since they are presented within the same interface and require zero or one step to revoke consent. According to Habib et al. (Habib et al., 2022), users who face persistent icons are more likely to recognise a correct method to revoke consent with respect to users who saw a link in the website’s footer. Therefore, even though all such implementations are compliant, only 7.4% websites with an Icon provide a more usable revocation design.
5.1.2. Two or more steps to revoke vs zero steps to accept
35 (22.1%) websites out of 158 allow users to access the revocation option within the same interface, but with additional obstruction, requiring them 2 or more steps to revoke consent. Such websites (labeled with Options via Policy) hid the option to revoke consent inside the “Cookie Policy” or “Privacy Policy” page. According to the majority of EU regulators, this additional effort does not allow to exercise the right of revocation in a easy and effective way even if such option is located withing the interface. Arguably, the websites requiring additional effort neither comply with LR3 (Easy revocation through the same effort and number of steps) nor with the principles P1 (Fairness), since websites require unjustifiable and unexpected effort, nor P2 (Data Protection by Design), since the adopted measures are not efficient as to facilitate the revocation right but obstruct it instead, and P3 (Accountability), since such websites are not able to demonstrate compliance with the requirement LR3.
5.1.3. Revocation options Via Different Interface
Overall, 32 (19.87%) websites out of 158 offered the option to revoke consent via an interface that is substantially different from the interface for users to accept consent. Figure 7 in the Appendix shows examples of websites for each of the options given.
Twenty-five websites, including medium.com, discord.com and wikipedia.org, suggest revocation through browser settings by clearing cookies or by offering opt-out links for third-party tracking or advertisement domains (labeled as Settings or links). Three websites – github.io, archive.org and who.int – suggested users to contact or email them to revoke consent or delete the data (labeled Contact/email).
Settings or links as well as Contact/email revocation options infringe the legal requirement LR2 (Easy revocation through the same interface), since revoking consent is not made available through the same means or interface; and principles P1 and P2, since switching to these totally different interfaces requires unnecessary and disruptive effort that is obstructive and unexpected. This leaves users in an asymmetrical relationship between giving and revoking consent, which does not permit users to exercise their revocation right. Consequently, these websites cannot be accountable for demonstrating compliance with this right (P3).
On one website(tumblr.com) privacy settings lead the user to the login page and instruct them to revoke consent after logging-in, which may require account creation (labeled After login). While this option occurs through the same website, it directs the user to a different interface. As before, this option infringes the requirement LR2 since the option redirects users to an unrelated interface. It also infringes principles P1 and P2 as it forces users to take unexpected additional steps and effort – to subscribe, and thus to give more personal data – to revoke consent. This option impedes these websites to show accountable revocation (P3).
5.1.4. No revocation option provided
Nine (5.6%) websites out of 158 did not provide any means to revoke consent, though they also did not display any consent banner, indicating that these websites probably did not integrate online tracking. We further analysed these websites and found that 4 (2.48%) of them stored AA cookies making them non-complaint. These websites include ntp.org and un.org. All these 4 websites that use AA cookies infringe the legal requirement LR1 (Right to revoke consent) and the following principles: P1 since websites process data unjustifiably for AA purposes without a legal basis, or the knowledge or consent of users; P2 as it does not create measures to enable the exercise of the revocation right. Consequently, these websites cannot be accountable for demonstrating compliance with this right (P3).
Figure 3. Revocation methods on top 158 websites
5.2. Impact of CMPs on compliance
Figure 3 shows the prevalence of different revocation options on the websites with detected CMPs vis a vis websites where we did not detect any CMP. 43 (77%) out of 56 websites with detected CMPs provide a compliant implementation (zero or one step) to revoke consent, while only 32 (30.5%) websites out of 105 websites without detected CMP have compliant implementation, showing overall a higher compliance rate for websites with CMPs.
Regarding non-compliant implementation, 37% (39 out of 105) websites without CMP require two steps or more to revoke consent, while only 10.7% (6 out of 56) websites with CMP needed 2 or more steps for revocation. None of the websites with CMP denied a revocation option to users, though 9 out of 105 websites without CMP offered no revocation option. In conclusion, websites with detected CMPs tend to be more compliant with revocation requirements.
6. Effect of Revocation on AA Cookies
(a) Comparison of AA cookies upon
initial landing vs after revocation
(b) Comparison of AA cookies
after rejection vs after revocation
(c) Comparison of AA cookies
after acceptance vs after revocation
Figure 4. Change in the number of advertising and analytics (AA) cookies across different options
In this section, we address RQ2 and measure whether cookies requiring consent (Advertising and Analytics, or AA cookies) are deleted upon user’s consent revocation. As described in § 4.1.2, we recorded the cookies of websites in four stages – upon initial landing, after accepting consent, after rejecting optional cookies, and after revocation. Out of the 120 websites where revocation was possible within the same interface, we find AA cookies after revocation on 69 (57.5%) websites.
Figure 4 compares the change in the number of AA cookies on websites after revocation w.r.t. other stages. Surprisingly, the number of AA cookies increases on most websites after revocation w.r.t. initial landing and w.r.t. after rejection (see red bar in Figures 4(a) and 4(b)). Additionally, on the majority of analysed websites, the number of AA cookies after acceptance remains the same after revocation (see pink bars in Figure 4(c)). Websites that add or keep AA cookies after consent revocation violate LR4 (Revoking requires stopping data processing and deletion of consent-based data), P1 and P2. Such AA cookies are processed unexpectedly and contrary to user’s decisions, without a legal basis and are thus illegal  (European Parliament and Council of the European Union, 2016, Art. 6(1)(a)).
7. Validity and Consistency of Consent
In this section, we analyse how consent information is stored and shared behind a website interface, whether it is consistently stored across different storage and APIs and whether it is legally valid, thus answering RQ3. We analysed 281 websites with detected CMPs (DS2 and DS3 from Table 2) to check the validity and consistency of the consent string that these CMPs implement. Together with a legal expert co-author of the paper, we analyzed 11 purposes predefined within IAB Europe TCF v2.2 (IAB, 2025b) (see Table 9 in the Appendix). These purposes are largely consistent with those of the TCF v2.0 (Matte et al., 2020b). We determined that purposes 2-9 require consent and demand user action to be selected, thus such purposes cannot be enabled by default in TCString. Purposes 10-11 are exempted from this requirement and may be enabled by default. Purpose 1 is storage-based and enabled by default.
Positive consent: if the given TCString contains at least one of the purposes from 2-9, and at least one vendor in its vendor list, we consider it to contain a positive consent. Such consent is correct only within the Acceptance phase, where the user has actively selected purposes 2-9 requiring user action.
Negative consent: if only purposes 1, 10 or 11 are enabled in the TCString, we conclude that it contains negative consent because none of these purposes require any user action as per our legal analysis. Consequently, if such TCString is present upon initial landing, or after rejection or after revocation stage, we consider consent to be registered correctly.
For the consent strings extracted from OTAG variable, or from OptanonConsent cookie, we follow the OneTrust specification (OneTrust, [n. d.]) and extract purpose numbers from its “groups” parameters. Since we do not have a specification for the meaning of these purposes, we cannot analyse which ones require consent. We therefore assume that a consent string contains a negative consent if it matches the value observed upon the initial landing stage, and a positive consent if the value contains more purposes than at initial landing.
Storage method or API Number of websites Positive consent
Total using
the method
Rejection
possible
Revocation
possible
Initial
landing
After
rejection
After
revocation
__tcfapi 163 150 136 1 2 17 (15 not updated)
TCF Cookies 130 104 101 0 0 12 (11 not updated)
TCF localStorage 41 31 32 0 1 8 (7 not updated)
OTAG 176 164 152 0 10 22 (16 not updated)
OptanonConsent Cookies 131 130 130 0 0 13 (10 not updated)
Table 5. Websites where consent strings in browser storage and returned by APIs were positive in different phases.
7.1. Results
7.1.1. Positive consent upon initial landing, after rejection and after revocation
We first analyse if the TCStrings returned by __tcfapi can be considered as a “baseline” in our comparison, i.e., whether such TCStrings are valid. For this purpose, we analysed whether the __tcfapi returns a positive consent at three stages: Initial landing, Rejection and Revocation, as shown in Table 5.
After Rejection and Initial landing: Only two prior works (Matte et al., 2020a; Zhang et al., 2024) analysed whether positive consent is stored or returned by APIs upon Initial landing and After Rejection. Compared to previous works, we found very few cases where positive consent is present after rejection in websites implementing the TCF (only 3 websites out of 150). In contrast, 10 websites, out of 164, store positive consent after rejection with OTAG.
After Revocation: We detected multiple websites where a positive consent is present after revocation, which has not been studied in previous works. In the websites with the TCF, 17 (12.5%) out of 136 websites provide positive consent in the __tcfapi, 12 (11.9%) out of 101 websites store a positive consent in a cookie, and 8 (25%) out of 32 websites store a positive consent in a localStorage. In the websites using OneTrust, we observe 22 (14.5%) out of 152 websites with positive consent returned by OTAG , and 13 (10%) out of 130 websites with positive consent in the OptanonConsent cookie. Positive consent constitutes potential violations of LR5 (Correct Consent Registration) which entails that the registered consent must be identical to the user’s choice in the user interface; and of the principle P2 since websites did not implement technical measures and safeguards to assure that revocation is done efficiently.
7.1.2. Consent strings are often not updated after revocation
Since positive consent persisted post-revocation in many cases, we examined whether consent strings were properly updated (Table 5). Among websites using __tcfapi, 15 (9.3%) did not update the consent string, while 11 (10.9%) and 7 (21.9%) failed to update the TCString in cookies and localStorage, respectively. Notably, sourceforge.com lacked a banner, initially stored negative consent, but changed it to positive upon revocation. ft.com updated only the “consentScreen” parameter, keeping positive consent intact. For websites using OTAG, 16 (10.5%) out of 152 failed to update consent post-revocation. Additionally, six websites modified only some purposes while still retaining positive consent. Interestingly, cisco.com, opendns.com, and webex.com added one purpose to the consent string after revocation.
7.1.3. Inconsistency among consent strings between browser storage and APIs
Table 6 shows the number of websites where the TCString found in the cookies and localStorage is not consistent with the TCString returned by the __tcfapi, and where the OptanonConsent cookie does not match the consent value returned by the OTAG. Such mismatches result in either an incorrect storage of consent or incorrect functioning of the APIs, and result in wrong consent being returned to third-parties. 5 websites implementing the TCF returned different TCStrings from __tcfapi and the TCString stored in the cookie after revocation. In 2 of such websites, (freep.com and megaphone.com) the TCString in the cookie was not updated, and still contained a positive consent, while the TCString returned by the __tcfapi was properly updated into a negative consent. In the website aol.com, while the __tcfapi returned a positive consent, the value in the cookie stores negative consent.
In three websites (reuters.com, manchestereveningnews.co.uk and portfolio.com), the TCString returned by __tcfapi showed negative consent, while the TCString stored in the localStorage remained unchanged, i.e, it contained positive consent. On websites implementing OneTrust, we observed 4 mismatches between OTAG and OptanonConsent cookie value, where the cookies stored a negative consent and were shared on the network. This mismatch can lead to the incorrect consent being shared to third-parties.
These mismatches between consent strings across browser storage and APIs can be due to the fact that websites do not consistently update the storage and APIs when users revoke consent. Consequently, websites do not correctly register consent, as they are obligated to, thus, infringing the consent requirement LR5. Moreover, P2 is not also complied with since websites do not implement technical measures to assure that revocation is done efficiently.
Browser storage vs. API # websites
TCF cookie vs. __tcfapi 5 (100)
TCF localStorage vs. __tcfapi 3 (32)
OT cookie vs. OTAG 4 (130)
Table 6. Inconsistent consent strings across browser storages and APIs after revocation. The numbers in brackets show the number websites that used both the storage and the API.
7.1.4. Inconsistency between consent returned via __tcfapi and consent shared on the network
We compare consent strings from __tcfapi with the consent strings found in outgoing network requests and incoming responses for two stages: Acceptance and Revocation. We consider the cases where the TCString returns the correct consent, i.e, a negative consent in case of revocation.
We investigate inconsistencies in the consent strings shared to and received from the third-parties, which, in turn, helps identify the responsible party for such inconsistencies. Previous works have only partially analysed such inconsistency (Matte et al., 2020a, §VIII.A), where positive consent was found to be sent within a specific gdpr_consent URL parameter on websites where __tcfapi did not contain a positive consent string. We however observed that the consent can be shared via differently named URL parameter as well. Additionally, the consent string can be sent in the POST data as part of the request as well. We find different inconsistencies on 8 distinct websites out of the 136 websites where revocation is possible.
These inconsistencies could arise due to: (1) delay in updating the consent after user revokes consent; (2) some scripts not being updated about revocation in consent; (3) introducing/using a different TCString. We do not consider delay in sending the updated consent as a violation since it is implementation specific. However, there are 8 websites where we observe inconsistencies causing possible violations. Out of these 8 websites, 4 of them (forbes.com, time.com, n-tv.de and cadenaser.com) had a different TCString on the network while deadline.com, kotaku.com, manchestereveningnews.co.uk and walesonline.co.uk, we observed that the old positive consent string was shared on the network by some scripts even after revocation. Details regarding these inconsistencies can be found in Appendix A.
The detected mismatches between consent strings returned by __tcfapi and the TCString shared on the network (Table 8) lead to an incorrect registration of consent revocation by websites and also by the implemented CMPs. Consequently, websites do not comply with LR5 since they do not correctly register user revocation, as they are obligated to, in order to assure that the registered consent is identical to the user’s choice in the user interface. As a result, a single TCString does not serve as a proof of revocation, which should consist of a negative consent. Moreover, the principle P2 is also not complied with since websites did not implement technical measures to assure that revocation is done efficiently.
8. Communicating Consent Revocation to Third-parties
Next, we address RQ4 and investigate if all the third-parties that were informed of the consent acceptance (either by accessing APIs or via HTTP requests) are also informed of consent revocation.
8.1. Results
8.1.1. Not all third-parties are informed of revocation by accessing consent via APIs
To examine which third-parties access the APIs implemented by CMPs, we override the implementation of __tcfapi and OTAG (see §4.2.4). Overall, we found 23 (9.6%) websites out of 238 where at least one third-party accessed the API to fetch positive consent after acceptance, but did not access the API to fetch revoked consent. Among these, on 163 websites where __tcfapi is implemented (see Table 5), on 14 (8.6%) websites at least one third-party requests the __tcfapi with the command “getTCData” after acceptance but not after revocation [^3] In the latest version of TCF, there is an option for third-parties to add an event-listener for consent update. However, since the implementation of this API is not standard, we could not intercept calls to it.. Out of 176 websites that support OTAG, on 13 (7.4%) websites at least one third-party accesses the OTAG variable after accepting, but not after revocation. On 4 of these 13 websites, scripts from the third-party domain ads-static.conde.digital read the value of the variable, indicating that this variable can be used to know the consent choice given for AA purposes, since this domain is present in EasyList (Fanboy and Khrin, 2023).
CMPs expose APIs to third-parties to access consent information making them responsible for requesting access to consent and for being informed on updates. Therefore, a proper implementation of event listeners by the website or by CMPs is necessary to inform third-parties about consent revocation, since consent can be updated multiple times in a single user session. We discuss about the standardization of the event-listeners in §9.
% 3rd-parties not informed Number of websites
<25absent25<25< 25% 1
[…]
100100100100% 45
Table 7. Percentage of third-parties informed of acceptance but not informed of the revocation in 101 websites.
8.1.2. Not all third-parties are informed of the consent revocation via HTTP requests
On each website where consent revocation is possible, we detected third-parties that are informed of consent information via the TCStrings and One Trust consent string in the URL or postData (see §4.2.4). We then detect third-parties informed of consent after acceptance but not informed of revocation.
Of the 136 websites that implement __tcfapi and support revocation, 101 (74.2%) websites contain at least one third-party that was informed of consent after acceptance, but not after revocation via HTTP requests. Moreover, on 68 of these websites, third-parties that were not informed about revocation set cookies through the “Set-Cookie” HTTP header, and thus, did not stop processing user’s data after revocation. Table 7 shows the percentage of third-parties not informed of consent after revocation. Surprisingly, 45 (44.5%) websites did not inform any of the detected third-parties, while 35 (34.6%) of them did not inform more than 75% of included third-parties. Figure 5 shows most prevalent third-parties that were not informed of the revoked consent via HTTP requests but were informed of consent acceptance. These websites include the big ad-tech industry actors doubleclick.net, criteo.com and adnxs.com. Interestingly, 73.77% of all third-party domains that do not receive a communication about consent revocation are present in EasyList (Fanboy and Khrin, 2023) and therefore participate in advertising or tracking.
Of the 152 websites that provide revocation and implement OneTrust, we observed that allowed categories from the OptanonConsent cookie (see §4.2.5) are rarely shared with third-parties, i.e., only on 8 websites OneTrust allowed categories were sent to third-parties after acceptance, and 6 sent the allowed categories after revocation. Interestingly, all third-party requests with allowed categories were made to Google Analytics (www.google-analytics.com) and Google Tag Manager (gtm.elementor.com). According to the official OneTrust documentation, the categories are communicated to the Google Tag Manager for it to manage consent and third parties (OneTrust, 2024)). We therefore conclude that OneTrust CMP may delegate the consent communication to the GTM framework and not send the allowed categories to third-parties by itself.
Figure 5. Top 15 third-parties not informed of the revoked consent via HTTP requests and the number of websites including these scripts but not informing them of revocation.
8.1.3. Failure to communicate consent revocation results in unlawful data processing
Our findings indicate that third-parties are not informed about consent revocation. This fact entails that these tracking- and advertising-based third-parties may continue to unlawfully process user’s personal data even after users have revoked consent on the website’s interface, since they were not updated of the user’s decision by the website. Consequently, websites from both cases (communication of the TCString via API and HTTP requests) are in potential violation with LR6 (Communication of withdrawal to third-parties) that demands websites to communicate consent revocation to the third-parties. The continuous unexpected processing of user’s data after revocation infringes the fairness principle P1 and the data protection by design principle P2 due to the lack of technical measures and safeguards efficiently enabling the exercise of the revocation right by websites.
9. Recommendations
In this section, we also propose recommendations to regulators for enhancing the implementation of consent revocation and ensuring legal compliance, in the light of our findings.
R1. Need to unify EU requirements for consent revocation interface. Our research shows that websites implement revocation inconsistently across the Web (see Table 4). For example, the text placed in the website footer differs across websites (e.g., “Cookie Settings”, “Privacy Settings”, but also “EU Privacy”), and in many cases does not mention revocation. It is even harder to locate revocation links within privacy policies that are labelled with text such as “How do I control cookies and how my data is used?”, or “Managing our analytics cookies”. The vague and ambiguous representation of these options misleads the users and prevents them from exercising the revocation right which should be clearly and distinctly recognisable (see LR1). Icons, that are supported by regulators and are easy to locate, use different visualisations, which may also confuse the user. EU DPAs should propose unified interface requirements describing the interface, location, and wordings of revocation. An inspiration can be taken from CPRA regulation in California, where three acceptable wordings are proposed, which simplifies locating it with automated means to measure compliance (Tran et al., 2024).
R2. Need to standardize consent storage. We observed various inconsistencies in storing the TCStrings and OneTrust consent strings (see §7). This is due to the usage of different storage options that are not updated simultaneously when consent is revoked. We invite regulators to establish standards for consent storage, and implement security measures to protect the integrity of consent strings.
R3. Need to standardize consent communication via event listeners. Our results show that only 43 out of 163 websites had third-parties installing event listeners in order to be updated about the change in consent decisions (see §8.1.1). While OneTrust suggests developers to add event listeners to synchronize consent updates when integrating advertising systems (OneTrust, 2024), our results show that IAB Europe TCF does not standardize event listeners, prompting every CMP to propose their own solution. This makes it harder for third-parties to adapt to different implementations across websites, placing the burden on them along with legal consequences (for continuous processing of personal data), without offering a practical solution to ensure compliance. We therefore recommend standard-setting bodies like IAB Europe TCF to standardize the implementation of event listeners and the callback functions returned by event listeners to inform third-parties about consent revocation. We also propose regulators to take a position on the means of communicating consent to third-parties. We believe that event listeners are able to provide reasonable means for communicating revocation decisions, and help in the allocation of responsibility to collect valid consent by third-parties present on a website.
R4. Need to regulate consent communication via HTTP requests. The most prevalent violation we detected indicates that 74.2% of websites do not properly inform all third-parties about consent revocation when the HTTP request method is used (see §8.1.1). Our result shows that this concrete way of informing third-parties exempts big advertising actors, such as doubleclick.net and criteo.com (see Figure 5) from any responsibility regarding data deletion when users revoke consent. Moreover, these third-parties only receive positive consent from the majority of websites and consequently, while illegally processing data from users who revoked consent, these companies can falsely demonstrate evidence of compliance to regulators by providing a false proof of consent. We therefore propose that EU regulators express their position regarding how websites should inform third-parties, and whether HTTP requests is an acceptable method, since our findings demonstrate that this approach leads to the most prevalent violations.
10. Conclusion
In this work, we propose a framework to audit compliance of consent revocation on the Web. We found multiple instances of violations of the EU Data Protection law while analysing interfaces of revocation, including use of different interfaces (19.87%) and more effort to revoke than to accept (20.5%). Multiple violations were observed on the usage of cookies and storing positive consent despite user’s revocation. Most shockingly, on 74% of websites, third-parties that have received consent upon user’s acceptance, are not informed of user’s revocation, leading to the illegal processing of users’ data by such third-parties. Our findings emphasise the need for improved legal compliance of consent revocation, in particular, proper, consistent, and uniform implementation of revocation communication and data deletion practices.
Acknowledgements
This work has been supported by the ANR 22-PECY-0002 IPoP (Interdisciplinary Project on Privacy) project of the Cybersecurity PEPR, the TULIP project of the ANR MRSEI program 2023, and the Inria International Chair funding.
Appendix
Appendix A Inconsistency between consent returned via __tcfapi and consent shared on the network
Responsible party Website Reason for inconsistency
3rd-party (CMP) Ketch Kloud, Inc. forbes.com Different TCString: number of vendor consents and number of vendor legitimate purposes increased in TCString on the network w.r.t to the TCString returned by the __tcfapi.
3rd-party (CMP) Ketch Kloud, Inc. time.com Different TCString: number of vendor consents, number of vendor legitimate purposes and number purpose consent increased in TCString on the network w.r.t TCString returned by the __tcfapi. 13 different TCStrings received in the responses.
1st-party n-tv.de Hardcoded TCString: String hardcoded in 1st-party script, sent in the request to 3rd-party. The TCString had additional purpose legitimate interests (5) and vendor legitimate interests (3).
3rd-party (CMP) Didomi cadenaser.com Third-party sends a different TCString w.r.t TCString returned by TCString in the response to a request made by a TP which was informed of the correct (TCString returned by __tcfapi) in previous requests.
1st-party deadline.com Old, non-updated (after acceptance) consent is shared via network requests even after revocation.
3rd-party (kinja-static.com) kotaku.com Third-party script was loaded after the consent was updated (from the previous stages, both in case of after acceptance and revocation), but it still used the wrong consent.
3rd-party (CMP) InMobi PTE Ltd manchester-eveningnews.co.uk, walesonline.co.uk Although all the requests sent an updated consent after revocation, signal-beacon.s-onetag.com/beacon.min.js script continues to send the positive consent (old consent-after acceptance)
Table 8. Inconsistencies in consent strings shared over the network
Purpose Name Repeated from v2.0 Requires Consent
1 Store and/or access information on a device = 1
2 Use limited data to select advertising Looks new but related to no. 2 = “Select basic ads”
3 Create profiles for personalised advertising = 3
4 Use profiles to select personalised advertising = 4
5 Create profiles to personalise content = 5 (✓) Not clear: in principle, this purpose can be legitimized under a legitimate interest, but it would fail the this test.
6 Use profiles to select personalised content = 6 (✓) Not clear: in principle, this purpose can be legitimized under a legitimate interest, but it would fail the legitimate interest test.
7 Measure advertising performance = 7
8 Measure content performance = 8
9 Understand audiences through statistics or combinations of data from different sources New language but looks like the former purpose no. 9, "Apply market research to generate audience insights"
10 Develop and improve services = 10 It is not specific, and so we cannot derive its legal basis.
X but in principle it could rely on LI though it could fail the LI test
11 Use limited data to select content New X
Table 9. IAB TCF purposes in v2.2 in the TCString and the applicable legal basis. The “Requires Consent” column sums up our analysis. The “Repeated from v2.0.” column compares the purposes of versions v2.0 and v2.2. We add parentheses if exceptions occur. You can find the descriptions of all the purposes in the supplementary material (Anonymous, 2025).
Table 8 presents different types of inconsistencies we have observed in our analysis, grouped by the reason for such inconsistency. We observed three types of inconsistencies in the requests and responses: (1) due to delay in updating the consent after user revokes consent; (2) due to introducing/using a different TCString; (3) due to some scripts not being updated about revocation in consent.
Cases where the consent was not updated immediately are the requests sent to third-parties and containing the TCString before user update. We also investigate the time difference between the user update event and when the old consent was still being sent on the network. We additionally observe cases where some URL parameters were updated while some variable in the POST data containing the old consent and taking few seconds to be updated. Examples include cnn.com, idnes.cz,independent.co.uk.
Cases where a completely different TCString was shared on the network, include two websites forbes.com and time.com, where the CMP’s server sends a TCString in the response to a request made by the CMP’s script. This TCString is not the same as the one returned by the __tcfapi and contains additional vendors and purposes. We however don’t observe this TCString being shared to other third-parties. We observe this behavior both after acceptance and after revocation. In case of time.com, we observed 13 different TCStrings. We also observe one case where a completely different TCString with additional legitimate interests in n-tv.de w.r.t the TCString retuned by the __tcfapi after both in case of after acceptance and after revocation. In this case, the TCString was hard-coded in a first-party script and sent to a third-party. In case of cadenaser.com, a third-party server sends a different TCString in response to a third-party script request. However it was not further sent to other third-parties. We also want to point out here that we observe four cases where the ConsentScreen parameter in the TCString is changed. However it does not affect the user’s consent in any way, hence is not relevant for the discussion.
Finally, we also observe cases where the scripts do not use the updated consent. While in deadline.com, all the requests use the old consent, in two websites, manchestereveningnews.co.uk and walesonline.co.uk some scripts send requests with the updated consent while the one script (signal-beacon.s-onetag.com/beacon.min.js) still sends the old consent on the network. Upon further investigation, we found that this script is used by the CMP but is not the one responsible for updating consent when the user takes action to do so. Table 8 also shows that localStorage was not updated while the cookies and __tcfapi was. It’s possible that the script used the value from local storage instead, resulting in an outdated consent being sent. However, there is no evidence to support this.
(a) Persistent icon
(b) Footer options
(c) Banner on the policy page
(d) Options on other pages
Figure 6. Revocation options within the interface on websites
(a) Browser settings or 3rd-party links
(b) Consent options after login
(c) Contact/Email
(d) Paywalls
(e) Revocation mentioned but not possible
(f) Privacy settings without links
Figure 7. Websites with revocation options via different interfaces or no revocation options
Within the Same Interface Via Different Interfaces [HTML]EA9999No Revocation
Banner No. of Websites Icon or Button (0 Steps) Footer Options (1 Step) Banner on Policy (1 step) Options through Policy Page ($\geq$ 2 steps) Browser set., 3rd-party links After Login Contact/ Email Paywalls Option mentioned but doesn’t work No revocation
Consent Banner 54 4 27 3 18 2 0 0 0 0 0
No Option Banner 8 0 0 0 5 2 1 0 0 0 0
No Banner 43 0 2 5 7 17 0 3 0 0 4 [+5]
Total 105 4 29 8 30 21 1 3 0 0 9
Table 10. Prevalence and type of consent revocation options on 105 websites (rank 1-200) without detected CMPs
Within the Same Interface Via Different Interfaces [HTML]EA9999No Revocation
Banner No. of Websites Icon or Button (0 Steps) Footer Options (1 Step) Banner on Policy (1 step) Options through Policy Page ($\geq$ 2 steps) Browser set., 3rd-party links After Login Contact/ Email Paywalls Option mentioned but doesn’t work No revocation
Consent Banner 54 5 36 3 3 6 1 0 0 0 0
No Option Banner 0 0 0 0 5 2 1 0 0 0 0
No Banner 2 0 0 0 0 0 0 0 0 0 0
Total 56 5 38 3 3 6 1 0 0 0 0
Table 11. Prevalence and type of consent revocation options on 56 websites (rank 1-200) with detected CMPs
Within the Same Interface Via Different Interfaces [HTML]EA9999No Revocation
Banner No. of Websites Icon or Button (0 Steps) Footer Options (1 Step) Banner on Policy (1 step) Options through Policy Page ($\geq$ 2 steps) Browser set., 3rd-party links After Login Contact/ Email Paywalls Option mentioned but doesn’t work No revocation
Consent Banner 214 33 125 0 26 9 0 3 11 3 7
No Option Banner 0 0 0 0 0 0 0 0 0 0 0
No Banner 11 0 8 0 0 1 0 0 0 0 2
Total 225 33 133 0 26 10 0 3 11 3 9
Table 12. Prevalence and type of consent revocation options on 200 to 5000 ranked websites with detected CMPs